### v1

# Policy Mirror Descent with Lookahead

Abstract
Policy Mirror Descent (PMD) stands as a versatile algorithmic framework encompassing several seminal policy gradient algorithms such as natural policy gradient, with connections with state-of-the-art reinforcement learning (RL) algorithms such as TRPO and PPO. PMD can be seen as a soft Policy Iteration algorithm implementing regularized 1-step greedy policy improvement. However, 1-step greedy policies might not be the best choice and recent remarkable empirical successes in RL such as AlphaGo and AlphaZero have demonstrated that greedy approaches with respect to multiple steps outperform their 1-step counterpart. In this work, we propose a new class of PMD algorithms called $h$-PMD which incorporates multi-step greedy policy improvement with lookahead depth $h$ to the PMD update rule. To solve discounted infinite horizon Markov Decision Processes with discount factor $\gamma$, we show that $h$-PMD which generalizes the standard PMD enjoys a faster dimension-free $\gamma^{h}$-linear convergence rate, contingent on the computation of multi-step greedy policies. We propose an inexact version of $h$-PMD where lookahead action values are estimated. Under a generative model, we establish a sample complexity for $h$-PMD which improves over prior work. Finally, we extend our result to linear function approximation to scale to large state spaces. Under suitable assumptions, our sample complexity only involves dependence on the dimension of the feature map space instead of the state space size.
1 Introduction
Policy Mirror Descent (PMD) is a general class of algorithms for solving reinforcement learning (RL) problems. Motivated by the surge of interest in understanding popular Policy Gradient (PG) methods, PMD has been recently investigated in a line of works (Lan, 2023; Xiao, 2022; Li et al., 2023; Johnson et al., 2023). Notably, PMD encompasses several PG methods as particular cases via its flexible mirror map. A prominent example is the celebrated Natural PG method. PMD has also close connections to state-of-the-art methods such as TRPO and PPO (Tomar et al., 2022) which have achieved widespread empirical success (Schulman et al., 2017a, b), including most recently in fine-tuning Large Language Models via RL from human feedback (Ouyang et al., 2022). Interestingly, PMD has also been inspired by one of the most fundamental algorithms in RL: Policy Iteration (PI). While PI alternates between policy evaluation and policy improvement, PMD regularizes the latter step to address the instability issue of PI with inexact policy evaluation.
Policy Iteration and its variants have been extensively studied in the literature, see e.g. Bertsekas and Tsitsiklis (1996); Munos (2003); Scherrer (2013). In particular, PI has been studied in conjunction with the lookahead mechanism (Efroni et al., 2018a), i.e. using multi-step greedy policy improvement instead of single-step greedy policies. Intuitively, the idea is that the application of the Bellman operator multiple times before computing a greedy policy leads to a more accurate approximation of the optimal value function. Implemented via Monte Carlo Tree Search (MCTS), multi-step greedy policy improvement [^1] Note that this is different from $n$-step return approaches which are rather used for policy evaluation (Sutton and Barto, 2018). contributed to the empirical success of some of the most impressive applications of RL including AlphaGo, AlphaZero and MuZero (Silver et al., 2016, 2017, 2018; Schrittwieser et al., 2019). Besides this practical success, a body of work (Efroni et al., 2018a, b; Tomar et al., 2020; Winnicki and Srikant, 2023; Winnicki et al., 2021) has investigated the role of lookahead in improving the performance of RL algorithms, reporting a convergence rate speed-up when enhancing PI with $h$-step greedy policy improvement with a reasonable computational overhead.
In this work, we propose to cross-fertilize the PMD class of algorithms with multi-step greedy policy improvement to obtain a novel class of PMD algorithms enjoying the benefits of lookahead. Our contributions are as follows:
  • We propose a novel class of algorithms called $h$-PMD enhancing PMD with multi-step greedy policy updates where $h$ is the depth of the lookahead. This class collapses to standard PMD when $h=1\,.$ When the step size parameter goes to infinity, we recover the PI algorithm with multiple-step greedy policy improvement previously analyzed in Efroni et al. (2018a). When solving a Markov Decision Process problem with $h$-PMD, we show that the value suboptimality gap converges to zero geometrically with a contraction factor of $\gamma^{h}$. This rate is faster than the standard convergence rate of PI and the similar rate for PMD which was recently established in Johnson et al. (2023) for $h\geq 2$. This rate improvement requires the computation of a $h$-step lookahead value function at each iteration which can be performed using planning methods such as tree-search. These results for exact $h$-PMD are exposed in Sec. 4, when value functions can be evaluated exactly.
  • We examine the inexact setting where the $h$-step action value functions are not available. In this setting, we propose a Monte Carlo sample-based procedure to estimate the unknown $h$-step lookahead action-value function at each state-action pair.
    We provide a sample complexity for this Monte Carlo procedure, improving over standard PMD thanks to the use of lookahead. Larger lookahead depth translates into a better sample complexity and a faster suboptimality gap convergence rate. However, this improvement comes at a more intensive computational effort to perform the lookahead steps. Sec. 5 discusses this inexact setting and the aforementioned tradeoff.
  • We extend our results to the function approximation setting to address the case of large state spaces where tabular methods are not tractable. In particular, we design a $h$-PMD algorithm where the $h$-step lookahead action-value function is approximated using a linear combination of state-action feature vectors in the policy evaluation step. Under linear function approximation and using a generative sampling model, we provide a performance bound for our $h$-PMD algorithm with linear function approximation. Our resulting sampling complexity only depends on the dimension of the feature map space instead of the size of the state space and improves over prior work analyzing PMD with function approximation.
  • We perform simulations to verify our theoretical findings empirically on the standard DeepSea RL environment from Deep Mind’s bsuite (Osband et al., 2020). Our experiments illustrate the convergence rate improvement of $h$-PMD with increasing lookahead depth $h$ in both the exact and inexact settings.
2 Preliminaries
Notation. For any integer $n\in\mathbb{N}\setminus\{0\}$, the set of integers from $1$ to $n$ is denoted by $[n]$. For a finite set $\mathcal{X},$ we denote by $\Delta(\mathcal{X})$ the probability simplex over the set $\mathcal{X}\,.$ For any integer $d\in\mathbb{N}$, we endow the Euclidean space ${{\mathbb{R}}}^{d}$ with the standard maximum norm $\|\cdot\|_{\infty}$ defined for every $v\in{{\mathbb{R}}}^{d}$ by $\|v\|_{\infty}:=\max_{i\in[d]}|v_{i}|$ where $v_{i}$ is the $i$-th coordinate. The relative interior of a set $\mathcal{Z}$ is denoted by $\operatorname{ri}(\mathcal{Z})\,.$
Markov Decision Process (MDP). We consider an infinite horizon discounted Markov Decision Process $\mathcal{M}=(\mathcal{S},\mathcal{A},r,\gamma,P,\rho)$ where $\mathcal{S}$ and $\mathcal{A}$ are finite sets of states and actions respectively with cardinalities $S=|\mathcal{S}|$ and $A=|\mathcal{A}|$ respectively, $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ is the Markovian probability transition kernel, $r:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is the reward function, $\gamma\in(0,1)$ is the discount factor and $\rho\in\Delta(\mathcal{S})$ is the initial state distribution. A randomized stationary policy is a mapping $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ specifying the probability $\pi(a|s)$ of selecting action $a$ in state $s$ for any $s\in\mathcal{S},a\in\mathcal{A}.$ The set of all such policies is denoted by $\Pi$. At each time step, the learning agent takes an action $a$ with probability $\pi(a|s)$, the environment transitions from the current state $s\in\mathcal{S}$ to the next state $s^{\prime}\in\mathcal{S}$ with probability $P(s^{\prime}|s,a)$ and the agent receives a reward $r(s,a)$.
Value functions and optimal policy. For any policy $\pi\in\Pi$, we define the state value function $V^{\pi}:\mathcal{S}\to{{\mathbb{R}}}$ for every $s\in\mathcal{S}$ by $V^{\pi}(s)=\operatorname{\mathbb{E}}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right]\,.$ The state-action value function $Q^{\pi}:\mathcal{S}\times\mathcal{A}\to{{\mathbb{R}}}$ can be similarly defined for every $s\in\mathcal{S},a\in\mathcal{A}$ by $Q^{\pi}(s,a)=\operatorname{\mathbb{E}}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a\right]$ where $\operatorname{\mathbb{E}}_{\pi}$ is the expectation over the state-action Markov chain $(s_{t},a_{t})$ induced by the MDP and the policy $\pi$ generating actions. The goal is to find a policy $\pi$ maximizing $V^{\pi}$. A classic result shows that there exists an optimal deterministic policy $\pi^{\star}\in\Pi$ maximizing $V^{\pi}$ simultaneously for all the states (Puterman, 2014). In this work, we focus on searching for an $\epsilon$-optimal policy, i.e. a policy $\pi$ satisfying $\|V^{\star}-V^{\pi}\|_{\infty}\leq\epsilon\,$ where $V^{\star}:=V^{\pi^{\star}}\,.$
Bellman operators and Bellman equations. We will often represent the reward function $r$ by a vector in ${{\mathbb{R}}}^{SA}$ and the transition kernel by the operator $P\in{{\mathbb{R}}}^{SA\times S}$ acting on vectors $v\in{{\mathbb{R}}}^{S}$ (which can also be seen as functions defined over $\mathcal{S}$) as follows: $(Pv)(s,a)=\sum_{s^{\prime}\in\mathcal{S}}P(s^{\prime}|s,a)v(s^{\prime})$ for any $s\in\mathcal{S},a\in\mathcal{A}.$ We further define the mean operator $M^{\pi}:{{\mathbb{R}}}^{SA}\to{{\mathbb{R}}}^{S}$ mapping any vector $Q\in{{\mathbb{R}}}^{SA}$ to a vector in ${{\mathbb{R}}}^{S}$ whose components are given by $(M^{\pi}Q)(s)=\sum_{a\in\mathcal{A}}\pi(a|s)Q(s,a)\,,$ and the maximum operator $M^{\star}:{{\mathbb{R}}}^{SA}\to{{\mathbb{R}}}^{S}$ defined by $(M^{\star}Q)(s)=\max_{a\in\mathcal{A}}Q(s,a)$ for any $s\in\mathcal{S},Q\in{{\mathbb{R}}}^{SA}\,.$ Using these notations, we introduce for any given policy $\pi\in\Pi$ the associated expected Bellman operator ${\mathcal{T}^{\pi}}:{{\mathbb{R}}}^{S}\to{{\mathbb{R}}}^{S}$ defined for every $V\in{{\mathbb{R}}}^{S}$ by ${\mathcal{T}^{\pi}}V:=M^{\pi}(r+\gamma PV)$ and the Bellman optimality operator ${\mathcal{T}}:{{\mathbb{R}}}^{S}\to{{\mathbb{R}}}^{S}$ defined by ${\mathcal{T}}V:=M^{\star}(r+\gamma PV)$ for every $V\in{{\mathbb{R}}}^{S}\,.$ Using the notations $r^{\pi}:=M^{\pi}r$ and $P^{\pi}:=M^{\pi}P$, we can also simply write ${\mathcal{T}^{\pi}}V=r^{\pi}+\gamma P^{\pi}V$ and ${\mathcal{T}}V=\max_{\pi\in\Pi}{\mathcal{T}^{\pi}}V$ for any $V\in{{\mathbb{R}}}^{S}\,.$
The Bellman optimality operator ${\mathcal{T}}$ and the expected Bellman operator ${\mathcal{T}^{\pi}}$ for any given policy $\pi\in\Pi$ are $\gamma$-contraction mappings w.r.t. the $\|\cdot\|_{\infty}$ norm and hence admit unique fixed points $V^{\star}$ and $V^{\pi}$ respectively. In particular, these operators satisfy the following inequalities: $\lVert{\mathcal{T}^{\pi}}V-V^{\pi}\rVert_{\infty}\leq\gamma\lVert V-V^{\pi}\rVert_{\infty}$ and $\lVert{\mathcal{T}}V-V^{\star}\rVert_{\infty}\leq\gamma\lVert V-V^{\star}\rVert_{\infty}$ for any $V\in{{\mathbb{R}}}^{S}\,.$ Moreover, we recall that for $\mathbf{e}=(1,1,\ldots,1)\in\mathbb{R}^{S}$ and any $c\in\mathbb{R}_{\geq 0},V\in{{\mathbb{R}}}^{S}$, we have ${\mathcal{T}^{\pi}}(V+ce)={\mathcal{T}^{\pi}}V+\gamma c\,\mathbf{e}$ and ${\mathcal{T}}(V+ce)={\mathcal{T}}V+\gamma c\,\mathbf{e}\,.$ We recall that the set $\mathcal{G}(V^{\star}):=\{\pi\in\Pi:{\mathcal{T}^{\pi}}V^{\star}={\mathcal{T}}V^{\star}\}$ of $1$step greedy policies w.r.t. the optimal value function $V^{\star}$ coincides with the set of stationary optimal policies.
3 A Refresher on PMD and PI with lookahead
3.1 PMD and its connection to PI
Policy Gradient methods consist in performing gradient ascent updates with respect to an objective $\mathbb{E}_{\rho}[V^{\pi}(s)]$ w.r.t. the policy $\pi$ parametrized by its values $\pi(a|s)$ for $s,a\in\mathcal{S}\times\mathcal{A}$ in the case of a tabular policy parametrization. Interestingly, policy gradient methods have been recently shown to be closely connected to a class of Policy Mirror Descent algorithms by using dynamical reweighting of the Bregman divergence with discounted visitation distribution coefficients in a classical mirror descent update rule using policy gradients. We refer the reader to section 4 in Xiao (2022) or section 3.1 in Johnson et al. (2023) for a detailed exposition.
The resulting PMD update rule is given by
$\pi^{k+1}_{s}\in\operatorname{argmax}_{\pi_{s}\in\Delta(\mathcal{A})}\left\{\eta_{k}\langle Q^{\pi_{k}}_{s},\pi_{s}\rangle-D_{\phi}(\pi_{s},\pi^{k}_{s})\right\}\,,$ (PMD)
where we use the shorthand notations $\pi^{k}_{s}={\pi_{k}}(\cdot|s)$ and $\pi_{s}=\pi(\cdot|s)$, where $(\eta_{k})$ is a sequence of positive step sizes, $Q^{\pi_{k}}_{s}\in\mathbb{R}^{A}$ is a vector containing state action values for ${\pi_{k}}$ at state $s$, and $D_{\phi}$ is the Bregman divergence induced by a mirror map $\phi:\operatorname{dom}\phi\to{{\mathbb{R}}}$ such that $\Pi\subset\operatorname{dom}\phi\,,$ i.e. for any $\pi,\pi^{\prime}\in\Pi,$
$D_{\phi}(\pi,\pi^{\prime})=\phi(\pi)-\phi(\pi^{\prime})-\langle\nabla\phi(\pi^{\prime}),\pi-\pi^{\prime}\rangle\,,$
where we suppose throughout this paper that the function $\phi$ is of Legendre type, i.e. strictly convex and essentially smooth in the relative interior of $\operatorname{dom}\phi$ (see Rockafellar (1970), section 26). The mirror map choice gives rise to a large class of algorithms. Two special cases are noteworthy: (a) When $\phi$ is the squared $2$-norm, the Bregman divergence is squared Euclidean distance and the resulting algorithm is a projected $Q$-ascent and (b) when $\phi$ is the negative entropy, the corresponding Bregman divergence is the Kullback-Leibler (KL) divergence and (PMD) becomes the popular Natural Policy Gradient algorithm (see e.g. Xiao (2022), section 4 for details).
Now, using our operator notations, one can observe that the (PMD) update rule above is equivalent to the following update rule (see Lemma A.1 for a proof),
$\pi_{k+1}\in\operatorname{argmax}_{\pi\in\Pi}\left\{\eta_{k}{\mathcal{T}^{\pi}}{V^{\pi_{k}}}-D_{\phi}(\pi,{\pi_{k}})\right\}\,,$ (1)
where $D_{\phi}(\pi,\pi_{k})\in\mathbb{R}^{S}$ is a vector whose components are given by $D_{\phi}(\pi_{s},\pi^{k}_{s})$ for $s\in\mathcal{S}\,.$ Note that the minimization can be carried out independently for each state component.
Connection to Policy Iteration. As previously noticed in the literature (Xiao, 2022; Johnson et al., 2023), PMD can be interpreted as a ‘soft’ PI. Indeed, taking infinite step sizes $\eta_{k}$ in PMD (or a null Bregman divergence) immediately leads to the following update rule:
$\pi_{k+1}\in\operatorname{argmax}_{\pi\in\Pi}\left\{{\mathcal{T}^{\pi}}{V^{\pi_{k}}}\right\}\,,$ (PI)
which corresponds to synchronous Policy Iteration (PI). The method alternates between a policy evaluation step to estimate the value function ${V^{\pi_{k}}}$ of the current policy $\pi_{k}$ and a policy improvement step implemented as a one-step greedy policy with respect to the current value function estimate.
3.2 Policy Iteration with Lookahead
As discussed in Efroni et al. (2018a), Policy Iteration can be generalized by performing $h\geq 1$ greedy policy improvement steps at each iteration instead of a single step. The resulting $h$-PI update rule can be written:
${\pi_{k+1}}\in\operatorname{argmax}_{\pi\in\Pi}\left\{{\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}\right\}\,,$ ($h$-PI)
where $h\in\mathbb{N}\setminus\{0\}\,.$ The $h$-greedy policy $\pi^{k+1}$ w.r.t $V^{\pi_{k}}$ selects the first optimal action of a non-stationary $h$-horizon optimal control problem. It can also be seen as the $1$-step greedy policy w.r.t. ${\mathcal{T}}^{h-1}V^{\pi_{k}}\,.$ In the rest of this paper, we will denote by $\mathcal{G}_{h}(V)$ the set of $h$-step greedy policies w.r.t. any value function $V\in{{\mathbb{R}}}^{S}$, i.e. $\mathcal{G}_{h}(V):=\operatorname{argmax}_{\pi\in\Pi}{\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}V=\{\pi\in\Pi:{\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}V={\mathcal{T}}^{h}V\}$. Notice that $\mathcal{G}_{1}(V)=\mathcal{G}(V)$. Overall, the $h$-PI method alternates between finding a $h$-greedy policy and estimating the value of this policy. Interestingly, similarly to PI, a $h$-greedy policy guarantees monotonic improvement, i.e. $V^{\pi^{\prime}}\geq V^{\pi}$ component-wise for any $\pi\in\Pi,\pi^{\prime}\in\mathcal{G}_{h}(V^{\pi})\,.$ Moreover, since ${\mathcal{T}}^{h}$ is a $\gamma^{h}$-contraction, it can be shown that the sequence $(\|V^{\star}-V^{\pi_{k}}\|_{\infty})_{k}$ is contracting with coefficient $\gamma^{h}\,.$ We refer the reader to section 3 in Efroni et al. (2018a) for a more detailed discussion regarding $h$-PI.
Remark 3.1.
While the iteration complexity always improves with larger depth $h$ (see Theorem 3 in Efroni et al. (2018a)), each iteration becomes more computationally demanding. As mentioned in Efroni et al. (2019), when the model is known, the $h$-greedy policy can be computed with Dynamic Programming (DP) in linear time in the depth $h$. When the model is unknown, one can use model-free algorithms such as Q-learning. Another possibility is to implement a tree-search of depth $h$ starting from the root state $s$ to compute the $h$-greedy policy in deterministic MDPs. We refer to Efroni et al. (2019) for further details. Sampling-based tree search methods such as Monte Carlo Tree Search (MCTS) (see Browne et al. (2012)) can be used to circumvent the exponential complexity in the depth $h$ of tree search methods.
4 Policy Mirror Descent with Lookahead
In this section we propose our novel Policy Mirror Descent (PMD) algorithm which incorporates $h$-step lookahead for policy improvement: $h$-PMD. Interestingly, this algorithm recovers the standard Policy Mirror Descent (PMD) (Xiao, 2022; Li et al., 2023) when $h=1$. Before introducing our algorithm, we start by describing the standard PMD algorithm and its connection with PI. Then we briefly discuss how multi-step lookahead has been incorporated to PI to motivate and present our novel $h$-PMD algorithm.
4.1 $h$-PMD: Using $h$-Step Greedy Updates
Similarly to the generalization of PI to $h$-PI discussed in the previous section, we obtain our algorithm $h$-PMD by incorporating lookahead to PMD as follows:
${\pi_{k+1}}=\operatorname{argmax}_{\pi\in\Pi}\left\{\eta_{k}{\mathcal{T}^{\pi}}{\mathcal{T}}^{h-1}{V^{\pi_{k}}}-D_{\phi}(\pi,{\pi_{k}})\right\}\,.$ ($h$-PMD)
Notice that we recover (1), i.e. (PMD), by setting $h=1\,.$ We now provide an alternative way to write the $h$-PMD update rule which is more convenient for its implementation. We introduce a few additional notations for this purpose. For any policy $\pi\in\Pi,$ let $V_{h}^{\pi}:={\mathcal{T}}^{h-1}V^{\pi}$ for any integer $h\geq~{}1\,.$ Consider the corresponding state-action value function defined for every $s,a\in\mathcal{S}\times\mathcal{A}$ by:
$Q_{h}^{\pi}(s,a):=r(s,a)+\gamma(PV_{h}^{\pi})(s,a)\,.$ (2)
Using these notations, it can be easily shown that the $h$-PMD rule above is equivalent to:
$\pi^{k+1}_{s}=\operatorname{argmax}_{\pi\in\Pi}\left\{\eta_{k}\langle Q_{h}^{\pi_{k}}(s,\cdot),\pi_{s}\rangle-D_{\phi}(\pi_{s},\pi_{s}^{k})\right\}\,.$ ($h$-PMD’)
Before moving to the analysis of this scheme, we provide two concrete examples for the implementation of $h$-PMD’ corresponding to the choice of two standard mirror maps.
(1) Projected $Q_{h}$-ascent. When the mirror map $\phi$ is the squared $2$-norm, the corresponding Bregman divergence is the squared Euclidean distance and $h$-PMD’ becomes
$\pi_{s}^{k+1}=\text{Proj}_{\Delta(\mathcal{A})}(\pi_{s}^{k}+\eta_{k}Q_{h}^{\pi_{k}}(s,\cdot))\,,$ (3)
for every $s\in\mathcal{S}$ and $\text{Proj}_{\Delta(\mathcal{A})}$ is the projection operator on the simplex $\Delta(A)\,.$
(2) Multiplicative $Q_{h}$-ascent. When the mirror map $\phi$ is the negative entropy, the Bregman divergence $D_{\phi}$ is the Kullback-Leibler divergence and for every $(s,a)\in\mathcal{S}\times\mathcal{A}$,
$\pi_{k+1}(a|s)=\pi_{k}(a|s)\frac{\exp(\eta_{k}Q_{h}^{\pi_{k}}(s,a)}{Z_{k}(s)}\,,$ (4)
where $Z_{k}(s):=\sum_{a\in\mathcal{A}}\pi_{k}(a|s)\exp(\eta_{k}Q_{h}^{\pi_{k}}(s,a))\,.$
4.2 Convergence Analysis for Exact $h$-PMD
Using the contraction properties of the Bellman operators, it can be shown that the suboptimality gap $\|V^{\pi_{k}}-V^{*}\|_{\infty}$ for PI iterates converges to zero at a linear rate with a contraction factor $\gamma$, regardless of the underlying MDP. This instance-independent convergence rate was generalized to PMD in Johnson et al. (2023). In this section, we establish a linear convergence rate for the subobtimality value function gap of the exact $h$-PMD algorithm with a contraction factor of $\gamma^{h}$ where $h$ is the lookahead depth. We assume for now that the value function $V^{\pi_{k}}$ and a greedy policy $\pi_{k+1}$ in ($h$-PMD) can be computed exactly. These assumptions will be relaxed in a subsequent theorem dealing with inexact $h$-PMD.
Theorem 4.1 (Exact $h$-PMD).
Let $(c_{k})$ be a sequence of positive reals and let the stepsize $\eta_{k}$ in ($h$-PMD) satisfy $\eta_{k}\geq\frac{1}{c_{k}}\lVert\min_{\pi\in\mathcal{G}_{h}({V^{\pi_{k}}})}D_{\phi}(\pi,\pi_{k})\rVert_{\infty}$ where we recall that $\mathcal{G}_{h}({V^{\pi_{k}}})$ is the set of greedy policies with respect to ${\mathcal{T}}^{h-1}{V^{\pi_{k}}}$ and that the minimum is computed component-wise. Initialized at $\pi_{0}\in\operatorname{ri}(\Pi)$, the iterates $(\pi_{k})$ of ($h$-PMD) with $h\in\mathbb{N}\setminus\{0\}$ satisfy for every $k\in\mathbb{N}$,
$\lVert V^{\star}-{V^{\pi_{k}}}\rVert_{\infty}\leq\gamma^{hk}\left(\lVert V^{\star}-V^{\pi_{0}}\rVert_{\infty}+\frac{1}{1-\gamma}\sum_{t=1}^{k}\frac{c_{t-1}}{\gamma^{ht}}\right).$
A few remarks are in order regarding this result:
  • Compared to Johnson et al. (2023), our new algorithm achieves a faster $\gamma^{h}$-rate where $h$ is the depth of the lookahead. It should be noted that the approach in Johnson et al. (2023) is claimed to be optimal over all methods that are guaranteed to increase the probability of the greedy action at each timestep. Adding lookahead to the policy improvement step circumvents this restriction.
  • Similarly to Johnson et al. (2023) and unlike prior work (Xiao, 2022; Lan, 2023) featuring distribution mismatch coefficients which can scale with the size of the state space, our convergence rate does not depend on any other instance-dependent quantities.
  • In Efroni et al. (2018a), it is shown that $h$-PI also achieves this same convergence rate, but without the additional term containing the errors due to $c_{k}$. This result can be recovered in $h$-PMD by setting the stepsize $\eta_{k}$ to infinity, effectively ignoring the effect of the divergence $D_{\phi}$. However, we have found experimentally that in the inexact case this leads to more instability, and a higher suboptimality gap at convergence.
  • As noted in Efroni et al. (2018a), this faster rate comes at the cost of a more complex value function computation at each iteration. However, our experiments show that the benefits of the faster convergence rate greatly outweigh the extra cost of computing the lookahead, in terms of both overall running time until convergence and sample complexity (in the inexact case).
  • The cost of computing the adaptive step sizes might seem high due to the minimisation over greedy policies. However as noted in Johnson et al. (2023) there is typically only one greedy policy at each iteration, and in the case that there is more than one it is possible to just pick any arbitrary greedy policy and the lower bound will still be satisfied.
  • We defer the proof of Theorem 4.1 to Appendix A. Our proof highlights the relationship between PMD and PI through the explicit use of Bellman operators. Notice that even in the particular case of $h=1$, our proof is more compact compared to the one in Johnson et al. (2023) (see sec. 6 and Lemma A.1 to A.3 therein for comparison) using our convenient notations.
5 Inexact $h$-Policy Mirror Descent
In this section, we discuss the case where the lookahead action value $Q_{h}^{\pi_{k}}$ defined in (2) is unknown. We propose a procedure to estimate it using Monte Carlo sampling and we briefly discuss an alternative using a tree search method. The resulting inexact $h$-PMD update rule for every $s\in\mathcal{S}$ is
$\pi_{s}^{k+1}=\operatorname{argmax}_{\pi_{s}\in\Delta(\mathcal{A})}\left\{\eta_{k}\langle\hat{Q}_{h}^{\pi_{k}}(s,\cdot),\pi_{s}\rangle-D_{\phi}(\pi_{s},\pi_{s}^{k})\right\},$ (5)
where $\hat{Q}_{h}^{\pi_{k}}$ is the estimated version of the lookahead action value $Q_{h}^{\pi_{k}}$ induced by policy $\pi_{k}\,.$ We conclude this section by providing a convergence analysis for inexact $h$-PMD and discussing its sample complexity using the proposed Monte Carlo estimator.
5.1 Monte Carlo $h$-Greedy Policy Evaluation
Estimating the lookahead action value function $Q_{h}^{\pi}$ for a given policy $\pi$ involves in particular solving a $h$-horizon planning problem using samples from the MDP. Our estimation procedure combines a planning method with Monte Carlo estimation under a generative model of the MDP.
We now describe our approach to estimate $Q_{h}^{\pi}(s,a)$ at a given fixed state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$. We define a sequence $(\mathcal{S}_{k})$ of sets of visited states as follows: Set $\mathcal{S}_{h}=\{s\}$ and define for every $k\in[0,h-1]$, for every $\tilde{s}\in\mathcal{S}_{k+1}$ and for ever $\tilde{a}\in\mathcal{A}$ the multiset $\mathcal{S}_{k}(\tilde{s},\tilde{a}):=\{s_{j}^{\prime}\sim P(\cdot|\tilde{s},\tilde{a}):j\in[M_{k+1}]\}$ (including duplicated states) where $M_{k+1}$ is the number of states sampled from each state $\tilde{s}\in\mathcal{S}_{k+1}\,,$ and $\mathcal{S}_{k}=\bigcup_{\tilde{s}\in\mathcal{S},\tilde{a}\in\mathcal{A}}\mathcal{S}_{k}(\tilde{s},\tilde{a})\cup\mathcal{S}_{k+1}$ (without duplicated states).
The estimation procedure is recursive. First we define the value function $\hat{V}_{1}$ as follows:
$\hat{V}_{1}(\tilde{s}):=\hat{V}^{\pi}(\tilde{s})\,,\quad\forall\tilde{s}\in\mathcal{S}_{0}\,,$ (6)
where $\hat{V}^{\pi}(\tilde{s})$ is an estimate of $V^{\pi}(\tilde{s})$ computed using $M_{0}$ rollouts of the policy $\pi$ from state $\tilde{s}\,.$ More precisely, we sample $M_{0}$ trajectories $(\tau_{j}(\tilde{s}))_{j\in[M_{0}]}$ of length $H\geq 1$ rolling out policy $\pi$, i.e. $\tau_{j}(\tilde{s}):=\{(s_{k}^{j},a_{k}^{j},r(s_{k}^{j},a_{k}^{j}))\}_{0\leq k\leq H-1}$ with $s_{0}^{j}=\tilde{s}$ and $a_{k}^{j}\sim\pi(\cdot|s_{k}^{j})$. Then the Monte Carlo estimator is computed as $\hat{V}^{\pi}(\tilde{s})=\frac{1}{M_{0}}\sum_{j=1}^{M_{0}}\sum_{k=0}^{H-1}\gamma^{k}r(s_{k}^{j},a_{k}^{j})\,.$ For each stage $k\in\left[1,h\right]$ for $h\geq 1$, we define the following estimates for every $\tilde{s}\in\mathcal{S}_{k},\tilde{a}\in\mathcal{A}$,
$\hat{Q}_{k}(\tilde{s},\tilde{a})$ $:=r(\tilde{s},\tilde{a})+\frac{\gamma}{M_{k}}\sum_{s^{\prime}\in\mathcal{S}_{k-1}}\hat{V}_{k}(s^{\prime})\,,$
$\hat{V}_{k+1}(\tilde{s})$ $:=\max_{\tilde{a}\in\mathcal{A}}\hat{Q}_{k}(\tilde{s},\tilde{a})\,,$ (7)
The desired estimate $\hat{Q}_{h}^{\pi}(s,a)$ is obtained with $k=h$. The recursive procedure described above can be represented as a (partial incomplete) tree where the root node is the state-action pair $(s,a)$ and each sampled action leads to a new state node $s^{\prime}$ followed by a new state-action node $(s^{\prime},a^{\prime})$. To trigger the recursion, we assign estimated values $\hat{V}^{\pi}(\tilde{s})$ of the true value function $V^{\pi}(\tilde{s})$ to each one of the leaf state nodes $\tilde{s}$ of this tree, i.e. the state level $h-1$ of this tree which corresponds to the last states visited in the sampled trajectories. The desired estimate $\hat{Q}_{h}^{\pi}(s,a)$ is given by the estimate obtained at the root state-action node $(s,a)\,.$ Recalling that $Q_{h}^{\pi}(s,a):=r(s,a)+\gamma(PV_{h}^{\pi})(s,a)$ for any $s\in\mathcal{S},a\in\mathcal{A}$ and $V_{h}^{\pi}:={\mathcal{T}}^{h-1}V^{\pi}$, the overall estimation procedure consists in using Monte Carlo rollouts to estimate the value function $V^{\pi}$ before approximating $V_{h}^{\pi}:={\mathcal{T}}^{h-1}V^{\pi}$ using approximate Bellman operator steps and finally computing the estimate $\hat{Q}_{h}^{\pi}(s,a)$ using a sampled version of $Q_{h}^{\pi}(s,a):=r(s,a)+\gamma(PV_{h}^{\pi})(s,a)$.
Remark 5.1.
Notice that actions are exhaustively selected in our procedure like in a planning method. Using bandit ideas to guide Monte Carlo planning (see e.g. Kocsis and Szepesvári (2006a)) would certainly be interesting to investigate for more efficiency. We leave the investigation of such selective action sampling and exploration for future work. In practice, lookahead policies are often computed using tree search techniques (Efroni et al., 2018a).
Remark 5.2.
When $h=1$, the procedure collapses to a simple Monte Carlo estimate of the Q-value function $Q^{\pi_{k}}$ at each iteration $k$ of the $h$-PMD algorithm.
5.2 Analysis of Inexact $h$-PMD
The next theorem is a generalization of Theorem 4.1 to the inexact setting in which the lookahead function $Q_{h}^{\pi_{k}}$ can be estimated with some errors at each iteration $k$ of $h$-PMD.
Theorem 5.3 (Inexact $h$-PMD).
Suppose there exists $b\in\mathbb{R}_{+}$ s.t. $\lVert\hat{Q}_{h}^{\pi_{k}}-Q_{h}^{\pi_{k}}\rVert_{\infty}\leq b$ where the maximum norm is over both the state and action spaces. Let $\tilde{\mathcal{G}}_{h}=\operatorname{argmax}_{\pi\in\Pi}\,\,M^{\pi}\hat{Q}^{\pi_{k}}_{h}$ and let $(c_{k})$ be a sequence of positive reals and let the stepsize $\eta_{k}$ in ($h$-PMD) satisfy $\eta_{k}\geq\frac{1}{c_{k}}\lVert\min_{\pi\in\tilde{\mathcal{G}}_{h}}D_{\phi}(\pi,\pi_{k})\rVert_{\infty}$ where we recall that the minimum is computed component-wise. Initialized at $\pi_{0}\in\operatorname{ri}(\Pi)$, the iterates $(\pi_{k})$ of inexact ($h$-PMD) with $h\in\mathbb{N}\setminus\{0\}$ satisfy for every $k\in\mathbb{N}$,
$\lVert V^{\star}-{V^{\pi_{k}}}\rVert_{\infty}\leq\gamma^{hk}\left(\lVert V^{\star}-V^{\pi_{0}}\rVert_{\infty}+\frac{1}{1-\gamma}\sum_{t=1}^{k}\frac{c_{t-1}}{\gamma^{ht}}\right)+\frac{2b}{(1-\gamma)(1-\gamma^{h})}\,.$
The proof of this result is similar to the proof of Theorem 4.1 and consists in propagating the error $b$ in the analysis. We defer it to Appendix B.
As can be directly seen from this formulation, the $\gamma^{h}$ convergence rate from the exact setting generalizes to the inexact setting. Thus, as before, faster rates of convergence are achieved for higher depths of lookahead. Another improvement with relation to Johnson et al. (2023) is that the additive error term has a weaker dependence on the effective horizon $1-\gamma$: the term $\frac{2b}{(1-\gamma)(1-\gamma^{h})}$ is smaller for all values of $h\geq 2$ than $\frac{4b}{(1-\gamma)^{2}}$. Therefore the algorithm converges to better solutions (in terms of suboptimality gap) as the asymptotic error on the suboptimality gap achievable by our algorithm (when $k\to\infty$) is smaller for deeper lookahead.
We are now ready to establish the sample complexity of inexact $h$-PMD under a generative model using Theorem 5.3 together with concentration results for our Monte Carlo lookahead action value estimator.
Theorem 5.4 (Sample complexity of $h$-PMD).
Assume that inexact $h$-PMD is run for a number of iterations $K>\frac{1}{h(1-\gamma)}\log(\frac{4}{\epsilon(1-\gamma)(1-\gamma^{h})})$, using the choice of stepsize defined by the sequence $(c_{k}):=(\gamma^{2h(k+1)})$. Additionally, suppose we are given a target suboptimality value $\epsilon>0$, and a probability threshold $\delta>0$. Finally, assume the lookahead value function $\hat{Q}_{h}^{\pi_{k}}$ is approximated at each iteration with the Monte Carlo estimation procedure described in section 5.1, with the following parameter values:
 $\forall\,1\leq j\leq h\,,M_{j}=M>\frac{144}{(1-\gamma)^{6}\varepsilon^{2}}\log(4h|\mathcal{A}||\mathcal{S}_{0}||\mathcal{C}|/\delta)\,,$
$M_{0}=\frac{144\gamma^{2h}}{(1-\gamma)^{4}(1-\gamma^{h})^{2}\varepsilon^{2}}\log(4|\mathcal{S}_{0}||\mathcal{C}|/\delta)\,,$
Then, with probability at least $1-\delta$, the suboptimality at all iterations $k$ satisfy the following bound:
$\lVert V^{\star}-{V^{\pi_{k}}}\rVert_{\infty}\leq\gamma^{hk}\left(\lVert V^{\star}-V^{\pi_{0}}\rVert_{\infty}+\frac{1-\gamma^{hk}}{(1-\gamma)(1-\gamma^{h})}\right)+\epsilon$ (8)
Using this procedure, $h$-PMD uses at most $KM_{0}H|\mathcal{S}|+hM|\mathcal{S}||\mathcal{A}|$ samples in total. The overall sample complexity of the inexact $h$-PMD is then given by $\tilde{\mathcal{O}}(\frac{|\mathcal{S}|}{h\epsilon^{2}(1-\gamma)^{6}(1-\gamma^{h})^{2}}+\frac{|\mathcal{S}||\mathcal{A}|}{\epsilon^{2}(1-\gamma)^{7}})$ where the notation $\tilde{O}$ hides at most polylogarithmic factors.
Compared to Theorem 5.1 in Johnson et al. (2023), the dependence on the effective horizon improves by a factor of the order of $1/(1-\gamma)$ for $h>|\mathcal{A}|^{-1}(1-\gamma)^{-1}$ thanks to our Monte Carlo lookahead estimator. Their sample complexity is of order $\tilde{\mathcal{O}}(\frac{1}{\epsilon^{2}(1-\gamma)^{8}})$, whereas for $h>\frac{1}{|\mathcal{A}|(1-\gamma)}$ ours becomes $\tilde{\mathcal{O}}(\frac{1}{\epsilon^{2}(1-\gamma)^{7}})$. Similarly to Johnson et al. (2023), our sample complexity does not depend on quantities such as the distribution mismatch coefficient used in prior work (Xiao, 2022; Lan, 2023), which may scale with the state space size. We provide a proof of this theorem including more detailed lower bounds for the parameters in the appendix.
6 Extension to Function Approximation
In MDPs with prohibitively large state action spaces, we represent each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ by a feature vector $\psi(s,a)\in{{\mathbb{R}}}^{d}$ where typically $d\ll|\mathcal{S}|\times|\mathcal{A}|$ and where $\psi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d}$ is a feature map, also represented as a matrix $\Psi\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times d}$. Assuming that this feature map holds enough “information” about each state action pair (more rigorous assumptions will be provided later), it is possible to approximate any value function using a matrix vector product $\Psi\theta$, where $\theta\in\mathbb{R}^{d}$. Following this idea, we propose a novel generalization of inexact $h$-PMD with function approximation by using this new value function for the inexact $h$-PMD update.
6.1 Inexact $h$-PMD with Function Approximation
Using the above notations, assume access to $\theta_{k}$, which we use to approximate $Q^{\pi_{k}}_{h}$ as $\Psi\theta_{k}$. Then the inexact $h$-PMD update rule can be adapted as follows:
$\pi_{s}^{k+1}\in\operatorname{argmax}_{\pi_{s}\in\Delta(\mathcal{S})}\left\{\eta_{k}\langle(\Psi\theta_{k})_{s},\pi_{s}\rangle-D_{\phi}(\pi_{s},\pi^{k}_{s})\right\}\,.$ (9)
We note that the policy update above can be implicit to avoid computing a policy for every state. It is possible to implement this algorithm using our Monte Carlo planning method in a similar fashion to Lattimore and Szepesvari (2019): use the procedure in 5.1 to estimate $\hat{Q}_{h}(s,a)$ for all state action pairs $(s,a)$ in some set $\mathcal{C}\subset\mathcal{S}\times\mathcal{A}$, and use these estimates as targets for a least squares regression to approximate $\hat{Q}_{h}$ by $\Psi\hat{\theta}$. We provide a specific procedure to accomplish this in the appendix. We conclude this section with a convergence analysis of $h$-PMD adapted to use linear function approximation.
6.2 Convergence Analysis of $h$-PMD with Linear Function Approximation
Our analysis follows the approach of Lattimore and Szepesvari (2019). We make the following standard assumptions on the feature map $\psi$, given by $\Psi$ in matrix form. We consider $h$ as a fixed input to the algorithm.
Assumption 6.1.
The feature matrix $\Psi$ is full rank.
Assumption 6.2 (Approximate Universal value function realizability).
There exists $\epsilon>0$ s.t. for any $\pi\in\Pi$, $\inf_{\theta\in\mathbb{R}^{d}}\lVert Q^{\pi}_{h}-\Psi\theta\rVert_{\infty}\leq\epsilon\,.$
Theorem 6.3 (Convergence of $h$-PMD with linear function approximation).
Let $(\theta_{k})$ be the sequence of iterates produced by $h$-PMD with linear function approximation, run for $K$ iterations, starting with policy $\pi_{0}\in\operatorname{ri}(\Pi)$, using stepsizes defined by $(c_{k})$. Assume that targets $\hat{Q}_{h}^{\pi_{k}}$ were computed using the procedure described in sec. 5.1 such that with probability at least $1-\delta$ for we have that $\forall k\;\leq K\,,\forall z\in\mathcal{C}\subset\mathcal{S}\times\mathcal{A}$ the targets satisfy $|\hat{Q}_{h}^{\pi_{k}}(z)-Q_{h}^{\pi_{k}}(z)|\leq\epsilon$ [^2] This can be achieved by choosing the parameters as in Theorem 5.4, with some slight modications presented in appendix C. Finally, for each iteration assume there exists some $\theta\in\mathbb{R}^{d}$ for which $\lVert Q^{\pi_{k}}_{h}-\Psi\theta\rVert_{\infty}\leq\lVert\epsilon_{\pi_{k}}\rVert_{\infty}$. There exists a choice of $\mathcal{C}$ for which the policies computed (or implicitly defined) at each iteration satisfy:
$\lVert V^{\star}-{V^{\pi_{k}}}\rVert_{\infty}$ $\leq\gamma^{hk}\left(\lVert V^{\star}-V^{\pi_{0}}\rVert_{\infty}+\frac{1}{1-\gamma}\sum_{t=1}^{k}\frac{c_{t-1}}{1-\gamma}\right)+\frac{2\max_{i\leq k}\lVert\epsilon_{\pi_{i}}\rVert_{\infty}(1+\sqrt{d})+2\sqrt{d}\,\epsilon}{(1-\gamma)(1-\gamma^{h})}\,.$
Notice that our performance bound does not depend on the state space size like in sec. 5 and depends instead on the dimension $d$ of the feature space. We should notice though that the choice of the set $\mathcal{C}$ is crucial for this result and given by the Kiefer-Wolfowitz theorem (Kiefer and Wolfowitz, 1960) which guarantees that $|\mathcal{C}|=\mathcal{O}(d^{2})$ as discussed in Lattimore and Szepesvari (2019).
As in the previous sections, our method for function approximation yields a sample complexity that is instance independent. In comparison to Alfano et al. (2023); Yuan et al. (2023), their results depend on distribution mismatch coefficients which can scale with the size of the state space.
Figure 1: Suboptimality value function gap for $h$-PMD in the exact (left) and inexact (right) settings. (Right) We performed 32 runs for each value of $h$, the mean is shown as a solid line and the standard deviation as a shaded area.
7 Simulations
We conduct simulations to investigate the effect of the lookahead depth on the convergence rate and illustrate our theoretical findings. We run the $h$-PMD algorithm for different values of $h$ in both exact and inexact settings on the DeepSea environment from DeepMind’s bsuite (Osband et al., 2020) using a grid size of 64 by 64, and a discount factor $\gamma=0.99$.
Exact $h$-PMD. We run the exact $h$-PMD algorithm for 100 iterations for increasing values of $h$ using the KL divergence. Similar results were observed for the Euclidean divergence.
We tested two different step size schedules: (a) in dotted lines in Fig. 1, $\eta_{k}$ equal to its lower bound in sec. 4, with the choice $c_{k}:=\gamma^{2h(k+1)}$ (note the dependence on $h$); and (b) in solid lines, $\eta_{k}$ identical stepsize schedule across all values of $h$ with $c_{k}:=\gamma^{2(k+1)}$ to isolate the effect of the lookahead. We clearly observe under both stepsize schedules that $h$-PMD converges in fewer iterations when lookahead value functions are used instead of regular $Q$ functions.
Inexact $h$-PMD. In this setting, we estimated the value function $Q_{h}$ using the vanilla Monte Carlo planning procedure detailed in section 5 in a stochastic variant of the DeepSea environment, and all the parameters except for $h$ were kept identical across runs. As predicted by our results, $h$-PMD converges in fewer iterations when $h$ is increased. We also observed in our simulations that $h$-PMD uses less samples overall (see Appendix D for the total number of samples used at each iteration), and usually converges after less overall computation time.
8 Related Work
Policy Mirror Descent and Policy Gradient Methods. Motivated by their empirical success (Schulman et al., 2017a, b), the analysis of PG methods has recently attracted a lot of attention (Agarwal et al., 2021; Bhandari and Russo, 2021; Khodadadian et al., 2022; Bhandari and Russo, 2024; Xiao, 2022; Lan, 2023; Johnson et al., 2023).
Among these works, a line of research focused on the analysis of PMD (Xiao, 2022; Lan, 2023; Li and Lan, 2023; Li et al., 2023; Johnson et al., 2023; Alfano et al., 2023) as a flexible algorithmic framework, and its particular instances such as the celebrated natural policy gradient (Khodadadian et al., 2022; Yuan et al., 2023). In particular, for solving unregularized MDPs (which is our main focus), Xiao (2022) established linear convergence of PMD in the tabular setting with a rate depending on the instance dependent distribution mismatch coefficients, and extended their results to the inexact setting using a generative model. Similar results were established for the particular case of the natural policy gradient algorithm (Bhandari and Russo, 2021; Khodadadian et al., 2022). More recently, Johnson et al. (2023) showed a dimension-free rate for PMD with exact policy evaluation using adaptive stepsizes, closing the gap with PI. Notably, they further show that this rate is optimal and their analysis inspired from PI circumvents the use of the performance difference lemma which was prevalent in prior work. While our analysis shares some similarity with this recent work in the tabular setting as we use similar adaptive stepsizes, we improve over these results using lookahead and we extend our analysis beyond the tabular setting by using linear function approximation. To the best of our knowledge, the use of multi-step greedy policy improvement in PMD and its analysis are novel in the literature.
Policy Iteration with Multiple-Step Policy Improvement. As mentioned, multi-step greedy policies have been succesfully used in applications such as the game of Go and have been studied in a body of works (Efroni et al., 2018a, b; Winnicki and Srikant, 2023; Efroni et al., 2020; Tomar et al., 2020). In particular, multi-step greedy policy improvement has been studied in conjunction with PI in Efroni et al. (2018a); Winnicki and Srikant (2023). (Efroni et al., 2018a) introduced $h$-PI which incorporates $h$-lookahead to PI and generalized existing analyses on PI with 1-step greedy policy improvement to $h$-step greedy policies. Our present work builds on these developments together with the connection between PMD and PI. Later, Efroni et al. (2018b) proposed softly and hardly updated multiple-step greedy policies for online and approximate PI. Inspired by this line of works, the class of $h$-PMD algorithms we introduce is novel to the best of our knowledge. Winnicki and Srikant (2023) showed that that a first-visit version of a PI scheme using a single sample path for policy evaluation converges to the optimal policy provided that the policy improvement step uses lookahead instead of a $1$-step greedy policy. Unlike our present work, it is assumed in the latter work that we have access to a lookahead greedy policy oracle.
Lookahead-Based Policy Improvement with Tree Search Lookahead policy improvement is usually implemented via tree search methods (Browne et al., 2012; Shah et al., 2020; Schrittwieser et al., 2019; Kocsis and Szepesvári, 2006b; Hallak et al., 2023; Efroni et al., 2019). Recently, Dalal et al. (2022) proposed a method combining policy gradient and tree search to reduce the variance of stochastic policy gradients. The method relies on a softmax policy parametrization incorporating a tree expansion. Our general $h$-PMD framework which incorporates a different lookahead policy improvement step also brings together PG and tree search methods. Morimura et al. (2022) designed a method with a mixture policy of PG and MCTS for non-Markov Decision Processes. Grill et al. (2020) showed that AlphaZero as well as several other MCTS algorithms compute approximate solutions to a family of regularized policy optimization problems which bear similarities with our $h$-PMD algorithm update rule for $h=1$ (see sec. 3.2 therein).
9 Conclusion
In this work, we introduced a novel class of PMD algorithms with lookahead inspired by the success of lookahead policies in practice as well as their analysis in PI with multi-step greedy improvements. We have shown an improved $\gamma^{h}$-linear convergence rate depending on the lookahead depth in the exact setting. We proposed an inexact version of our algorithm estimating the lookahead value function via planning and Monte Carlo sampling and we derived an improved sample complexity. We further extended our results to scale to large state spaces via linear function approximation with a performance bound independent on the state space size. Our paper offers several interesting directions for future research. A possible avenue for future work is to investigate the use of more general function approximators such as neural networks to approximate the lookahead value functions and scale to large state-action spaces. Recent work for PMD along these lines (Alfano et al., 2023) might be a good starting point. Enhancing our algorithm with exploration mechanisms (see e.g. the recent work O’Donoghue (2023)) is an important future direction. Our PMD update rule might offer some advantage in this regard as was recently observed in the literature (Li and Lan, 2023). Constructing and studying more efficient estimators for the lookahead action values using for example Monte Carlo Tree Search is also certainly a promising avenue for future work. Designing adaptive lookahead strategies to select the tree search horizon (Rosenberg et al., 2023) is yet another opportunity for further research. Our algorithm brings together two popular and successful families of RL algorithms: PG methods and tree search methods. Looking forward, we hope our work further stimulates research efforts to investigate practical and efficient methods enjoying the benefits of both classes of methods in solving practical RL problems.
Appendix A Additional Details and Proofs for Section 4: Exact $h$-PMD
Lemma A.1.
For any policies $\pi,\pi^{\prime}\in\Pi$ and every state $s\in\mathcal{S},$ we have $\langle Q^{\pi^{\prime}}_{s},\pi_{s}\rangle={\mathcal{T}^{\pi}}V^{\pi^{\prime}}(s)\,.$ As a consequence, the update rules (PMD) and (1) in the main part are equivalent.
Proof.
Let $\pi,\pi^{\prime}\in\Pi$. For every $s\in\mathcal{S}$, we have
$\langle Q^{\pi^{\prime}}_{s},\pi_{s}\rangle$ $=\sum_{a\in\mathcal{A}}Q^{\pi^{\prime}}(s,a)\,\pi(a|s)$ (by definition of the scalar product)
$=\sum_{a\in\mathcal{A}}\left(r(s,a)+\gamma\sum_{s^{\prime}\in\mathcal{S}}P(s^{\prime}|s,a)V^{\pi^{\prime}}(s^{\prime})\right)\pi(a|s)$ ($Q^{\pi^{\prime}}$ as a function of $V^{\pi^{\prime}}$)
$=r^{\pi}(s)+\gamma\sum_{s^{\prime}\in\mathcal{S}}\left(\sum_{a\in\mathcal{A}}P(s^{\prime}|s,a)\pi(a|s)\right)V^{\pi^{\prime}}(s^{\prime})$ (by rearranging and using the definition of $r^{\pi}$)
$=r^{\pi}(s)+\gamma P^{\pi}V^{\pi^{\prime}}(s)$ (using the definition of $P^{\pi}$)
$={\mathcal{T}^{\pi}}V^{\pi^{\prime}}(s)\,.$ (10)
The second part of the statement holds by setting $\pi^{\prime}=\pi_{k}\,.$
A.1 Auxiliary Lemma for the Proof of Theorem 4.1: Exact $h$-PMD
Lemma A.2 (Triple Point Descent Lemma for $h$-PMD).
Let $h\in\mathbb{N}\setminus\{0\}\,.$ Initialized at $\pi_{0}\in\operatorname{ri}(\Pi)$, the iterates of $h$-step greedy PMD remain in $\operatorname{ri}(\Pi)$ and satisfy the following inequality for any $k\in\mathbb{N}$ and any policy $\pi\in\Pi$,
$\eta_{k}{\mathcal{T}^{\pi}}{\mathcal{T}}^{h-1}{V^{\pi_{k}}}\leq\eta_{k}{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+D_{\phi}(\pi,{\pi_{k}})-D_{\phi}(\pi,{\pi_{k+1}})-D_{\phi}({\pi_{k+1}},{\pi_{k}})$
Proof.
Recall the definition of the $h$-step greedy action value function denoted by $Q^{\pi}_{h}:\mathcal{S}\times\mathcal{A}\to{{\mathbb{R}}}$ in (2). Then, recall that for each state $s\in\mathcal{S}$ and every $k\in\mathbb{N}$, we have
$\langle Q^{\pi_{k}}_{h}(s,\cdot),\pi_{s}\rangle=({\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}V^{\pi_{k}})(s)\,.$ (11)
Applying the three point descent lemma (see Lemma 6 in Xiao [2022] with $\phi(\pi)=\langle Q_{h}^{\pi_{k}}(s,\cdot),\pi\rangle$ for any $\pi\in\Pi$ and $\mathcal{C}=\Delta(\mathcal{A})$ with the notation therein), we obtain for any policy $\pi\in\Pi$,
$-\eta_{k}\langle Q^{\pi_{k}}_{d}(s,\cdot),\pi^{k+1}_{s}\rangle+D_{\phi}(\pi^{k+1}_{s},\pi^{k}_{s})\leq-\eta_{k}\langle Q^{\pi_{k}}_{d}(s,\cdot),\pi_{s}\rangle+D_{\phi}(\pi_{s},\pi^{k}_{s})-D_{\phi}(\pi_{s},\pi^{k+1}_{s})\,.$
We obtain the desired result by rearranging the above last inequality and using (11). ∎
Lemma A.3.
Let $h\in\mathbb{N}\setminus\{0\}\,.$ Initialized at $\pi_{0}\in\operatorname{ri}(\Pi)$, the iterates of $h$-step greedy PMD satisfy for every $k\in\mathbb{N},$
${\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}\leq{V^{\pi_{k+1}}}+\frac{\gamma}{1-\gamma}c_{k}\mathbf{e}\,,$
where $(c_{k})$ is the sequence of positive integers introduced in Theorem 4.1 such that $\frac{\tilde{d}_{k}}{\eta_{k}}\leq c_{k}\,$ and $\tilde{d}_{k}=\lVert\min_{\pi\in\mathcal{G}_{h}(V^{\pi_{k}})}D_{\phi}(\pi,{\pi_{k}})\rVert_{\infty}$.
Proof.
Let $k\in\mathbb{N}\,.$ Recall that any policy $\pi\in\mathcal{G}_{h}(V^{\pi_{k}})$ satisfies ${\mathcal{T}^{\pi}}{\mathcal{T}}^{h-1}{V^{\pi_{k}}}={\mathcal{T}}^{h}{V^{\pi_{k}}}\,.$ Using Lemma A.2 for $\pi\in\mathcal{G}_{h}(V^{\pi_{k}})$ together with the previous identity and the nonnegativity of the Bregman divergence implies that for any for any policy $\pi\in\mathcal{G}_{h}(V^{\pi_{k}})$,
${\mathcal{T}}^{h}{V^{\pi_{k}}}\leq{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+\frac{\tilde{d}_{k}}{\eta_{k}}\mathbf{e}\leq{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+c_{k}\mathbf{e}\,,$ (12)
where the last inequality follows from the assumption on the step sizes $\eta_{k}\,.$ We now prove by recurrence that for every $n\in\mathbb{N}$,
${\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}\leq({\mathcal{T}^{\pi_{k+1}}})^{n+1}{\mathcal{T}}^{h-1}{V^{\pi_{k}}}+\sum_{i=1}^{n}\gamma^{i}c_{k}\mathbf{e}\,,$ (13)
with $\sum_{i=1}^{n}\gamma^{i}=0$ for $n=0\,.$ The base case for $n=0$ is immediate. We now suppose that the result holds for $n$ and we show that it remains true for $n+1\,.$ We have
${\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}$ $={\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{\mathcal{T}}^{\pi_{k}}{V^{\pi_{k}}}$
$\leq{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}}^{h}{V^{\pi_{k}}}$
$\leq{\mathcal{T}^{\pi_{k+1}}}({\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+c_{k}\mathbf{e})$
$=({\mathcal{T}^{\pi_{k+1}}})^{2}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+\gamma c_{k}\mathbf{e}$
$\leq{\mathcal{T}^{\pi_{k+1}}}\left(({\mathcal{T}^{\pi_{k+1}}})^{n+1}{\mathcal{T}}^{h-1}{V^{\pi_{k}}}+\sum_{i=1}^{n}\gamma^{i}c_{k}\mathbf{e}\right)+\gamma c_{k}\mathbf{e}$
$=({\mathcal{T}^{\pi_{k+1}}})^{2}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+\sum_{i=1}^{n}\gamma^{i+1}c_{k}\mathbf{e}+\gamma c_{k}\mathbf{e}$
$=({\mathcal{T}^{\pi_{k+1}}})^{2}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+\sum_{i=1}^{n+1}\gamma^{i}c_{k}\mathbf{e}\,,$ (14)
which concludes the recurrence proof of (13) for every $n\in\mathbb{N}\,.$ To conclude, we take the limit in (13) and use the fact that for any $V\in{{\mathbb{R}}}^{S}$, the sequence $\left(({\mathcal{T}^{\pi_{k+1}}})^{n}V\right)_{n}$ converges to the fixed point ${V^{\pi_{k+1}}}$ of the contractive operator ${\mathcal{T}^{\pi_{k+1}}}\,.$ Hence, we obtain
${\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}\leq{V^{\pi_{k+1}}}+\sum_{i=1}^{+\infty}\gamma^{i}c_{k}\mathbf{e}={V^{\pi_{k+1}}}+\frac{\gamma}{1-\gamma}c_{k}\mathbf{e}\,.$ (15)
This concludes the proof. ∎
Remark A.4.
We note that although in our work the stepsise $\eta_{k}$ is chosen to be the same for all states for simplicity, the same result holds for a stepsize that is different for each state as long as $\eta_{k}(s)\geq\frac{1}{c_{k}}D_{\phi}(\pi_{s},\pi_{s}^{k})$ for each $s\in\mathcal{S}$. The same is true for all results depending on this lemma (or the inexact variant).
A.2 Proof of Theorem 4.1
As a warmup, we start by reporting a convergence result for the standard PMD algorithm, recently established in Johnson et al. [2023]. Using the contraction properties of the Bellman operators, it can be shown that the suboptimality gap $\|V^{\pi_{k}}-V^{*}\|_{\infty}$ for PI iterates converges to zero at a linear rate with a contraction factor $\gamma$, regardless of the underlying MDP. This instance-independent convergence rate was generalized to PMD in Johnson et al. [2023]. Before stating this result, we report a useful three-point descent lemma to quantify the improvement of the PMD policy resulting from the proximal step compared to an arbitrary policy.
Lemma A.5 (Three Point Descent Lemma).
Using the PMD update rule, we have for every $k\in\mathbb{N}$ and every $\pi\in\Pi$,
$-\eta_{k}{\mathcal{T}^{\pi_{k+1}}}{V^{\pi_{k}}}$ $\leq-\eta_{k}{\mathcal{T}^{\pi}}{V^{\pi_{k}}}-D_{\phi}({\pi_{k+1}},{\pi_{k}})+D_{\phi}(\pi,{\pi_{k}})-D_{\phi}(\pi,{\pi_{k+1}})\,.$ (16)
This lemma results from an application of Lemma 6 in Xiao [2022] which is a variant of Lemma 3.2 in Chen and Teboulle [1993]. This same result was also used in Johnson et al. [2023].
Theorem A.6 (Exact PMD, Johnson et al. [2023]).
Let $(c_{k})$ be a sequence of positive reals and let the stepsize $\eta_{k}$ in (PMD) satisfy $\eta_{k}\geq\frac{1}{c_{k}}\lVert\min_{\pi\in\mathcal{G}({V^{\pi_{k}}})}D_{\phi}(\pi,\pi_{k})\rVert_{\infty}$ where the minimum is computed component-wise. Initialized at $\pi_{0}\in\operatorname{ri}(\Pi)$, the iterates $(\pi_{k})$ of (PMD) satisfy for every $k\in\mathbb{N}$,
$\lVert V^{\star}-{V^{\pi_{k}}}\rVert_{\infty}\leq\gamma^{k}\left(\lVert V^{\star}-V^{\pi_{0}}\rVert_{\infty}+\sum_{t=1}^{k}\gamma^{-t}c_{t-1}\right)\,.$
We provide here a complete proof of this result as a warmup for our upcoming main result. Although equivalent, notice that our proof is more compact compared to the one provided in Johnson et al. [2023] (see section 6 and Lemma A.1 to A.3 therein) thanks to the use of our convenient notations, and further highlights the relationship between PMD and PI through the explicit use of Bellman operators.
Proof.
Using Lemma A.5 with a policy $\pi\in\mathcal{G}({V^{\pi_{k}}})$ for which ${\mathcal{T}^{\pi}}{V^{\pi_{k}}}={\mathcal{T}}{V^{\pi_{k}}}$, we have
$\eta_{k}{\mathcal{T}}{V^{\pi_{k}}}$ $\leq\eta_{k}{\mathcal{T}^{\pi_{k+1}}}{V^{\pi_{k}}}-D_{\phi}({\pi_{k+1}},{\pi_{k}})+D_{\phi}(\pi,{\pi_{k}})-D_{\phi}(\pi,{\pi_{k+1}})\,.$ (17)
Setting $d_{k}:=\lVert\min_{\pi\in\mathcal{G}({V^{\pi_{k}}})}D_{\phi}(\pi,\pi_{k})\rVert_{\infty}$ and recalling that $D_{\phi}(\pi,\pi_{k})\geq 0$ for any policy $\pi\in\Pi$, we obtain
${\mathcal{T}}{V^{\pi_{k}}}\leq{\mathcal{T}^{\pi_{k+1}}}{V^{\pi_{k}}}+\frac{d_{k}}{\eta_{k}}\mathbf{e}\,.$ (18)
Notice that for Policy Iteration we have ${\mathcal{T}}{V^{\pi_{k}}}={\mathcal{T}^{\pi_{k+1}}}{V^{\pi_{k}}}$ i.e. ${\pi_{k+1}}\in\mathcal{G}({V^{\pi_{k}}})$. Instead, for PMD we have that ${\pi_{k+1}}$ is approximately greedy (see (18)).
Applying Lemma A.5 again with $\pi=\pi_{k}$, recalling that ${\mathcal{T}^{\pi_{k}}}{V^{\pi_{k}}}={V^{\pi_{k}}}$ and that $D_{\phi}(\pi,\pi^{\prime})\geq 0$ for any $\pi,\pi^{\prime}\in\Pi$, we immediately obtain ${V^{\pi_{k}}}\leq{\mathcal{T}^{\pi_{k+1}}}{V^{\pi_{k}}}\,.$ Iterating this relation and using the monotonicity of the Bellman operator, we obtain ${V^{\pi_{k}}}\leq{\mathcal{T}^{\pi_{k+1}}}{V^{\pi_{k}}}\leq({\mathcal{T}^{\pi_{k+1}}})^{2}{V^{\pi_{k}}}\leq\cdots\leq({\mathcal{T}^{\pi_{k+1}}})^{n}{V^{\pi_{k}}}\leq\cdots\leq{V^{\pi_{k+1}}}\,$ where the last inequality follows from the fact that ${\mathcal{T}^{\pi_{k+1}}}$ has a unique fixed point ${V^{\pi_{k+1}}}\,.$ We are now ready to control the suboptimality gap as follows:
$V^{\star}-{V^{\pi_{k+1}}}$ $\leq V^{\star}-{\mathcal{T}^{\pi_{k+1}}}{V^{\pi_{k}}}$
$=V^{\star}-{\mathcal{T}}{V^{\pi_{k}}}+{\mathcal{T}}{V^{\pi_{k}}}-{\mathcal{T}^{\pi_{k+1}}}{V^{\pi_{k}}}$
$\leq V^{\star}-{\mathcal{T}}{V^{\pi_{k}}}+\frac{d_{k}}{\eta_{k}}\mathbf{e}$
$\leq V^{\star}-{\mathcal{T}}{V^{\pi_{k}}}+c_{k}\,,$
where the first inequality follows from the inequality ${V^{\pi_{k+1}}}\geq{\mathcal{T}^{\pi_{k+1}}}{V^{\pi_{k}}}$ proved above, the second one stems from (18), and the last one is a consequence of our assumption $d_{k}/\eta_{k}\leq c_{k}\,.$ We conclude the proof by taking the max norm, applying the triangle inequality and using the contractiveness of the Bellman optimality operator ${\mathcal{T}}\,.$
Proof of Theorem 4.1
Proof.
Similarly to Lemma A.5, we show that for any $\pi\in\mathcal{G}_{h}({V^{\pi_{k}}})$,
$\eta_{k}{\mathcal{T}^{\pi}}{\mathcal{T}}^{h-1}{V^{\pi_{k}}}$ $\leq\eta_{k}{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}-D_{\phi}({\pi_{k+1}},{\pi_{k}})+D_{\phi}(\pi,{\pi_{k}})-D_{\phi}(\pi,{\pi_{k+1}})\,,$ (19)
see Lemma A.2 for a proof. Setting $\tilde{d}_{k}:=\lVert\min_{\pi\in\mathcal{G}_{h}({V^{\pi_{k}}})}D_{\phi}(\pi,\pi_{k})\rVert_{\infty}$ (note the difference with $d_{k}$ in the proof of Theorem A.6), and recalling that $\pi\in\mathcal{G}_{h}({V^{\pi_{k}}})$ implies that $\pi\in\mathcal{G}({\mathcal{T}}^{h-1}{V^{\pi_{k}}})$, we obtain from (19) an inequality similar to (18),
${\mathcal{T}}^{h}{V^{\pi_{k}}}\leq{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}}^{h-1}{V^{\pi_{k}}}+\frac{\tilde{d}_{k}}{\eta_{k}}\mathbf{e}\,.$ (20)
Then, we observe that
${\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}$ $={\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{\mathcal{T}^{\pi_{k}}}{V^{\pi_{k}}}$
$\leq{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}}^{h}{V^{\pi_{k}}}$
$\leq{\mathcal{T}^{\pi_{k+1}}}({\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}}^{h-1}{V^{\pi_{k}}}+\frac{\tilde{d}_{k}}{\eta_{k}}\mathbf{e})$
$=({\mathcal{T}^{\pi_{k+1}}})^{2}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+\gamma\frac{\tilde{d}_{k}}{\eta_{k}}\mathbf{e}$
$\leq({\mathcal{T}^{\pi_{k+1}}})^{2}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+\gamma c_{k}\mathbf{e}\,,$ (21)
where the second inequality follows from using (19) and monotonicity of ${\mathcal{T}^{\pi_{k+1}}}$ and the last inequality uses the assumption on the step sizes. Then we show by recursion by iterating (A.2), taking the limit and using the fact that ${\mathcal{T}^{\pi_{k+1}}}$ has a unique fixed point ${V^{\pi_{k+1}}}$, that
${\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}\leq{V^{\pi_{k+1}}}+\frac{\gamma}{1-\gamma}c_{k}\mathbf{e}\,,$ (22)
see Lemma A.3 for a complete proof. We now control the suboptimality gap as follows,
$V^{\star}-{V^{\pi_{k+1}}}$ $\leq V^{\star}-{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+\frac{\gamma}{1-\gamma}c_{k}\mathbf{e}$
$=V^{\star}-{\mathcal{T}}^{h}{V^{\pi_{k}}}+\frac{\gamma}{1-\gamma}c_{k}\mathbf{e}+{\mathcal{T}}^{h}{V^{\pi_{k}}}-{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}$
$\leq V^{\star}-{\mathcal{T}}^{h}{V^{\pi_{k}}}+\frac{1}{1-\gamma}c_{k}\mathbf{e}\,,$ (23)
where the first inequality stems from (22) and the last one from reusing (20). Similarly to Theorem A.6, we conclude the proof by taking the max norm, applying the triangle inequality and using the contractiveness of the Bellman optimality operator $h$ times to obtain
$\lVert V^{\star}-{V^{\pi_{k+1}}}\rVert_{\infty}\leq\gamma^{h}\lVert V^{\star}-{V^{\pi_{k}}}\rVert_{\infty}+\frac{1}{1-\gamma}c_{k}\,.$
A recursion gives the desired result. ∎
Appendix B Additional Details and Proofs for Section 5: Inexact $h$-PMD
B.1 Proof of Theorem 5.3: Inexact $h$-PMD
Throughout this section, we suppose that the assumptions of Theorem 5.3 hold.
Lemma B.1.
Let $\pi\in\Pi,k\in\mathbb{N}\,.$ Suppose that the estimator $\hat{Q}^{{\pi_{k}}}_{h}$ is such that $\lVert\hat{Q}^{{\pi_{k}}}_{h}-Q^{\pi}_{h}\rVert_{\infty}\leq b$ for some positive scalar $b\,.$ Then the approximate $h$-step lookahead value function $\hat{V}_{h}^{\pi_{k}}(\pi):=M^{\pi}\hat{Q}^{\pi_{k}}_{h}$ is a $b$-approximation of ${\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}$, i.e.,
$\lVert{\hat{V}_{h}^{{\pi_{k}}}}(\pi)-{\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}\lVert_{\infty}\leq b\,.$ (24)
Proof.
The Hölder inequality implies that for any policy $\pi\in\Pi$:
$\lVert\hat{V}^{{\pi_{k}}}_{h}(\pi)-{\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}\rVert_{\infty}=\max_{s\in\mathcal{S}}|\langle\hat{Q}_{h}^{\pi_{k}}(s,\cdot)-Q_{h}^{\pi_{k}}(s,\cdot),\pi_{s}\rangle|\leq\max_{s\in\mathcal{S}}\lVert\hat{Q}_{h}^{\pi_{k}}(s,\cdot)-Q_{h}^{\pi_{k}}(s,\cdot)\rVert_{\infty}\cdot\lVert\pi_{s}\rVert_{1}\leq b\,.$
Lemma B.2 (Approximate $h$-Step Greedy Three-Point Descent Lemma).
For any policy $\pi\in\Pi$, the iterates $\pi_{k}$ of approximate $h$-PMD satisfy the following inequality for every $k\geq 0$:
$\eta_{k}{\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}\leq\eta_{k}{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+D_{\phi}(\pi,{\pi_{k}})-D_{\phi}(\pi,{\pi_{k+1}})-D_{\phi}({\pi_{k+1}},{\pi_{k}})+2\eta_{k}b\mathbf{e}\,.$ (25)
Proof.
We apply again the Three-Point Descent Lemma (see e.g. Lemma A.1 in Johnson et al. [2023]) to obtain:
$-\eta_{k}\langle{\hat{Q}_{h}^{\pi_{k}}}(s,\cdot),\pi_{s}^{k+1}\rangle+D_{\phi}(\pi_{s}^{k+1},\pi_{s}^{k})\leq-\eta_{k}\langle{\hat{Q}_{h}^{\pi_{k}}}(s,\cdot),\pi_{s}\rangle+D_{\phi}(\pi_{s},\pi_{s}^{k})-D_{\phi}(\pi_{s},\pi_{s}^{k+1})\,,$ (26)
which by rearranging and expressing in vector form gives:
$\eta_{k}{\hat{V}_{h}^{{\pi_{k}}}}(\pi)\leq\eta_{k}{\hat{V}_{h}^{{\pi_{k}}}}({\pi_{k+1}})+D_{\phi}(\pi,{\pi_{k}})-D_{\phi}(\pi,{\pi_{k+1}})-D_{\phi}({\pi_{k+1}},{\pi_{k}})\,.$ (27)
Lemma B.1 implies that ${\hat{V}_{h}^{{\pi_{k}}}}(\pi)\geq{\mathcal{T}^{\pi}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}-b\mathbf{e}$ and ${\hat{V}_{h}^{{\pi_{k}}}}({\pi_{k+1}})\leq{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+b\mathbf{e}$ yielding the lemma. ∎
The following lemma is an analogue of (20) in the proof of Theorem 4.1 above, modified for the inexact case.
Lemma B.3.
Assuming access at each iteration to an approximate $Q_{h}$ value function $\hat{Q}^{{\pi_{k}}}_{h}$ with $\lVert\hat{Q}^{{\pi_{k}}}_{h}-Q^{\pi}_{h}\rVert_{\infty}\leq b$, the iterates of inexact $h$-PMD satisfy for every $k\geq 0$:
${\mathcal{T}}^{h}{V^{\pi_{k}}}\leq{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+2b\mathbf{e}+\frac{\tilde{d}_{k}}{\eta_{k}}\mathbf{e}\,.$ (28)
Proof.
Let $\bar{\pi}$ be any true $h$-step greedy policy $\bar{\pi}\in\mathcal{G}_{h}({V^{\pi_{k}}})$, and let $\tilde{\pi}$ be any approximate $h$-step greedy policy $\tilde{\pi}\in\underset{\pi\in\Pi}{\operatorname{argmax}}\,\hat{V}_{h}^{\pi_{k}}(\pi)$. By definition of $\tilde{\pi}$ and equation (27) we have:
${\hat{V}_{h}^{{\pi_{k}}}}(\bar{\pi})\leq{\hat{V}_{h}^{{\pi_{k}}}}(\tilde{\pi})$ $\leq{\hat{V}_{h}^{{\pi_{k}}}}({\pi_{k+1}})+\frac{1}{\eta_{k}}D_{\phi}(\tilde{\pi},{\pi_{k}})-\frac{1}{\eta_{k}}D_{\phi}(\tilde{\pi},{\pi_{k+1}})-\frac{1}{\eta_{k}}D_{\phi}({\pi_{k+1}},{\pi_{k}})$ (29)
$\leq{\hat{V}_{h}^{{\pi_{k}}}}({\pi_{k+1}})+\frac{\tilde{d}_{k}}{\eta_{k}}\mathbf{e}\,.$ (30)
Finally we have the desired result using Lemma B.1, recalling the argument from the end of Lemma B.2 that ${\hat{V}_{h}^{{\pi_{k}}}}(\bar{\pi})\geq{\mathcal{T}}^{\bar{\pi}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}-b\mathbf{e}={\mathcal{T}}^{h}{V^{\pi_{k}}}-b\mathbf{e}$ and ${\hat{V}_{h}^{{\pi_{k}}}}({\pi_{k+1}})\leq{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+b\mathbf{e}$.
Lemma B.4.
Given a $\hat{Q}^{{\pi_{k}}}_{h}$ with $\lVert\hat{Q}^{{\pi_{k}}}_{h}-Q^{\pi}_{h}\rVert_{\infty}\leq b$, let $\tilde{\mathcal{G}}_{h}=\operatorname{argmax}_{\pi\in\Pi}{\hat{V}_{h}^{{\pi_{k}}}}(\pi)$ denote the set of approximate h-step greedy policies and $\tilde{d}_{k}=\lVert\min_{\tilde{\pi}\in\tilde{\mathcal{G}}_{h}}D(\tilde{\pi},{\pi_{k}})\rVert_{\infty}$. Then the iterates of inexact h-PMD satisfy the following bound:
$\eta_{k}{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}\leq\eta_{k}{V^{\pi_{k+1}}}+2\eta_{k}\frac{\gamma}{1-\gamma}b\mathbf{e}+\frac{\gamma}{1-\gamma}\tilde{d}_{k}\mathbf{e}\,.$ (31)
Proof.
Analogously to Lemma A.3 we proceed by induction using Lemma B.3. We write for every $k\geq 0$,
$\eta_{k}{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}$ $\stackrel{{\scriptstyle(a)}}{{\leq}}\eta_{k}{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}}^{h}{V^{\pi_{k}}}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\eta_{k}({\mathcal{T}^{\pi_{k+1}}})^{2}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+2\gamma\eta_{k}b\mathbf{e}+\gamma\tilde{d}_{k}\mathbf{e}$
$\leq\eta_{k}({\mathcal{T}^{\pi_{k+1}}})^{2}{\mathcal{T}}^{h}{V^{\pi_{k}}}+2\gamma\eta_{k}b\mathbf{e}+\gamma\tilde{d}_{k}\mathbf{e}$
$\leq\eta_{k}({\mathcal{T}^{\pi_{k+1}}})^{3}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+2\gamma^{2}\eta_{k}b\mathbf{e}+\gamma^{2}\tilde{d}_{k}+2\gamma\eta_{k}b\mathbf{e}+\gamma\tilde{d}_{k}\mathbf{e}$
[…]
$\leq\eta_{k}{V^{\pi_{k+1}}}+2\eta_{k}\frac{\gamma}{1-\gamma}b\mathbf{e}+\frac{\gamma}{1-\gamma}\tilde{d}_{k}\mathbf{e}\,,$ (32)
where (a) follows from the monotonicity of both Bellman operators ${\mathcal{T}^{\pi_{k+1}}},{\mathcal{T}}$ and the fact that ${V^{\pi_{k}}}\leq{\mathcal{T}}{V^{\pi_{k}}}$, (b) uses Lemma B.3 and monotonicity of the Bellman operators. The last inequality stems from the fact that for every $V\in{{\mathbb{R}}}^{S},\lim_{n\to+\infty}({\mathcal{T}^{\pi_{k+1}}})^{n}V={V^{\pi_{k+1}}}$ since ${V^{\pi_{k+1}}}$ is the unique fixed point of the Bellman operator ${\mathcal{T}^{\pi_{k+1}}}$ which is a $\gamma$-contraction. ∎
End of the proof of Theorem 5.3. We are now ready to conclude the proof of the theorem. Letting $\tilde{d}_{k}=\lVert\min_{\tilde{\pi}\in\tilde{\mathcal{G}}_{h}}D(\tilde{\pi},{\pi_{k}})\rVert_{\infty}$ as in the theorem, we have:
$V^{\star}-{V^{\pi_{k+1}}}$ $\stackrel{{\scriptstyle(i)}}{{\leq}}V^{\star}-{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+\frac{\gamma}{1-\gamma}\frac{\tilde{d}_{k}}{\eta_{k}}\mathbf{e}+2\frac{\gamma}{1-\gamma}b\mathbf{e}$
$=V^{\star}-{\mathcal{T}}^{h}{V^{\pi_{k}}}+{\mathcal{T}}^{h}{V^{\pi_{k}}}-{\mathcal{T}^{\pi_{k+1}}}{\mathcal{T}^{h-1}}{V^{\pi_{k}}}+\frac{\gamma}{1-\gamma}\frac{\tilde{d}_{k}}{\eta_{k}}\mathbf{e}+2\frac{\gamma}{1-\gamma}b\mathbf{e}$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}V^{\star}-{\mathcal{T}}^{h}{V^{\pi_{k}}}+\frac{1}{1-\gamma}\frac{\tilde{d}_{k}}{\eta_{k}}\mathbf{e}+\frac{2}{1-\gamma}b\mathbf{e}\,,$
where (i) follows from B.3 and (ii) from Lemma B.4. By the triangle inequality and the contraction property of ${\mathcal{T}}$ we obtain a recursion similar to the previous theorems:
$\lVert V^{\star}-{V^{\pi_{k+1}}}\rVert_{\infty}\leq\gamma^{h}\lVert V^{\star}-{V^{\pi_{k}}}\rVert_{\infty}+\frac{1}{1-\gamma}\frac{\tilde{d}_{k}}{\eta_{k}}+\frac{2}{1-\gamma}b\,.$
Unfolding this recursive inequality concludes the proof.
B.2 Proof of Theorem 5.4
We first recall the notations used in our lookahead value function estimation procedure before providing the proof. The procedure we use is a standard online planning method and our proof follows similar steps to the proof provided for example in the lecture notes [Szepesvári, 2023]. Notice though that we are not approximating the optimal Q function but rather the lookahead value function $Q^{\pi}_{h}$ induced by a given policy $\pi\in\Pi$ at any state action pair $(s,a)\in\mathcal{S}\times\mathcal{A}\,.$
Recall that for any $k\in\left[2,h-1\right],\hat{V}_{k}$ and $\hat{Q}_{k}$ are defined as follows:
$\hat{Q}_{k}(\tilde{s},\tilde{a})$ $=r(\tilde{s},\tilde{a})+\gamma\frac{1}{M_{k}}\sum_{s^{\prime}\in\mathcal{S}_{k-1}}\hat{V}_{k}(s^{\prime})\,,\quad\forall\tilde{s}\in\mathcal{S}_{k}\,,$ (33)
$\hat{V}_{k}(\tilde{s})$ $=\max_{a\in\mathcal{A}}\hat{Q}_{k-1}(\tilde{s},a)\,,\quad\forall\tilde{s}\in\mathcal{S}_{k-1}\,.$ (34)
At the bottom of the recursion, recall that $\hat{V}_{1}$ is defined for every $\tilde{s}\in\mathcal{S}_{0}$ by
$\hat{V}_{1}(\tilde{s})=\hat{V}^{\pi}(\tilde{s})=\frac{1}{M_{0}}\sum_{j=1}^{M_{0}}\sum_{k=0}^{H-1}\gamma^{k}r(s_{k}^{j},a_{k}^{j})\,,$
using $M_{0}$ trajectories $(\tau_{j}(\tilde{s}))_{j\in[M_{0}]}$ of length $H\geq 1$ rolling out policy $\pi$, i.e. $\tau_{j}(\tilde{s}):=\{(s_{k}^{j},a_{k}^{j},r(s_{k}^{j},a_{k}^{j}))\}_{0\leq k\leq H-1}$ with $s_{0}^{j}=\tilde{s}$ and $a_{k}^{j}\sim\pi(\cdot|s_{k}^{j})$.
Note that the cardinality of each set $\mathcal{S}_{k}$ is evidently upper bounded by $|\mathcal{S}|$ but also by $|\mathcal{A}|^{h-k}\prod_{i\in\left[k+1,h\right]}M_{i}$ as the number of states in each set grows by a factor of at most $|\mathcal{A}|M_{k}$ at step $k$. In this section we prove some concentration bounds on the above estimators, which will be useful for the rest of the analysis.
Lemma B.5 (Concentration Bounds of Lookahead Value Function Estimators).
Let $\delta\in(0,1)\,.$ For any state $\tilde{s}\in\mathcal{S}_{0}$, with probability $1-\delta$ we have:
$|\hat{V}_{1}(\tilde{s})-V^{\pi}(\tilde{s})|\leq C_{1}^{V}(\delta):=\frac{\gamma^{H}}{1-\gamma}+\frac{1}{1-\gamma}\sqrt{\frac{\log(2/\delta)}{M_{1}}}\,.$ (35)
For any $k\in\left[1,h\right]$, for state $\tilde{s}\in\mathcal{S}_{k}$ and for any action $\tilde{a}\in\mathcal{A}$, with probability $1-\delta$ we have:
$|\hat{Q}_{k}(\tilde{s},\tilde{a})-(J\hat{V}_{k})(\tilde{s},\tilde{a})|\leq C_{k}^{Q}(\delta):=\frac{\gamma}{1-\gamma}\sqrt{\frac{\log(2/\delta)}{M_{k}}}\,.$ (36)
Finally, for any $k\in\left[1,h-1\right]$, for any $\tilde{s}\in\mathcal{S}_{k}$, also with probability $1-\delta$ we have:
$|\hat{V}_{k+1}(\tilde{s})-({\mathcal{T}}\hat{V}_{k})(\tilde{s})|\leq C_{k+1}^{V}(\delta):=\frac{\gamma}{1-\gamma}\sqrt{\frac{\log(2|\mathcal{A}|/\delta)}{M_{k}}}\,.$ (37)
Proof.
The proof of this result is standard and relies on Hoeffding’s inequality. We provide a proof for completeness. Let $\Pr$ be the probability measure induced by the interaction of the planner and the MDP simulator on an adequate probability space. We denote by $\mathbb{E}$ the corresponding expectation.
First we show the bound for the value function estimate $\hat{V}_{1}$. This estimate is biased due to the truncation of the rollouts at a fixed horizon $H$, which is the source of the first error term in the bound. For any state $\tilde{s}\in\mathcal{S}_{0}$, the bias in $\hat{V}_{1}(\tilde{s})$ is given by
$|\operatorname{\mathbb{E}}\left[\hat{V}_{1}(\tilde{s})\right]-V^{\pi}(\tilde{s})|$ $=\operatorname{\mathbb{E}}\left[\sum_{t=H}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=\tilde{s},a_{t}\sim\pi(\cdot|s_{t})\right]\leq\frac{\gamma^{H}}{1-\gamma}\,.$ (38)
The second term is due to Hoeffding’s inequality: each trajectory is i.i.d, and the estimator is bounded with probability $1$ by $\frac{1}{1-\gamma}$. Therefore $\hat{V}_{1}(\tilde{s})$ concentrates around its mean, and with probability $1-\delta$ we have:
$|\hat{V}_{1}(\tilde{s})-\operatorname{\mathbb{E}}\left[\hat{V}_{1}(\tilde{s})\right]|\leq\frac{1}{1-\gamma}\sqrt{\frac{\log(2/\delta)}{M_{1}}}\,.$ (39)
Now we prove the bound for the $\hat{Q}_{k}$ function. Assume $k\in\left[1,h\right]$ is arbitrary, and $\tilde{s}\in\mathcal{S}_{k}$. For any action $\tilde{a}\in\mathcal{A}$, we can bound the error of $\hat{Q}_{k}$ as follows:
$|\hat{Q}_{k}(\tilde{s},\tilde{a})-(J\hat{V}_{k})(\tilde{s},\tilde{a})|=\gamma\left|\frac{1}{M_{k}}\sum_{i=1}^{M_{k}}\hat{V}_{k}(s^{\prime}_{i})-(P\hat{V}_{k})(\tilde{s},\tilde{a})\right|\leq\frac{\gamma}{1-\gamma}\sqrt{\frac{\log(2/\delta)}{M_{k}}}\,.$ (40)
Finally we bound the $\hat{V}_{k+1}$ function for $k>1$. Assume $k\in\left[1,h-1\right]$ is arbitrary, and $\tilde{s}\in\mathcal{S}_{k}$. With probability at least $1-\delta$ we have:
$|\hat{V}_{k+1}(\tilde{s})-({\mathcal{T}}\hat{V}_{k})(\tilde{s})|\leq\frac{\gamma}{1-\gamma}\sqrt{\frac{\log(2|\mathcal{A}|/\delta)}{M_{k}}}\,.$ (41)
To show this last result, we use a standard union bound. For every $t\geq 0$, we have:
$\Pr(|\hat{V}_{k+1}(\tilde{s})-{\mathcal{T}}\hat{V}_{k}(\tilde{s})|\geq t)$
$=\Pr(|\max_{a^{\prime}\in\mathcal{A}}\hat{Q}_{k}(\tilde{s},a^{\prime})-\max_{a^{\prime}\in\mathcal{A}}(JV_{k})(\tilde{s},a^{\prime})|\geq t)$
$\leq\Pr(\max_{a^{\prime}\in\mathcal{A}}\hat{Q}_{k}(\tilde{s},a^{\prime})-\max_{a^{\prime}\in\mathcal{A}}(JV_{k})(\tilde{s},a^{\prime})\geq t)+\Pr(\max_{a^{\prime}\in\mathcal{A}}(JV_{k})(\tilde{s},a^{\prime})-\max_{a^{\prime}\in\mathcal{A}}\hat{Q}_{k}(\tilde{s},a^{\prime})\geq t)$
$\leq\Pr(\max_{a^{\prime}\in\mathcal{A}}\left[\hat{Q}_{k}(\tilde{s},a^{\prime})-(JV_{k})(\tilde{s},a^{\prime})\right]\geq t)+\Pr(\max_{a^{\prime}\in\mathcal{A}}\left[(JV_{k})(\tilde{s},a^{\prime})-\hat{Q}_{k}(\tilde{s},a^{\prime})\right]\geq t)$
$\leq\sum_{\tilde{a}\in\mathcal{A}}\Pr(\hat{Q}_{k}(\tilde{s},\tilde{a})-(J\hat{V}_{k})(\tilde{s},\tilde{a})\geq t)+\sum_{\tilde{a}\in\mathcal{A}}\Pr((J\hat{V}_{k})(\tilde{s},\tilde{a})-\hat{Q}_{k}(\tilde{s},\tilde{a})\geq t)$
$\leq 2|\mathcal{A}|\exp(-2\frac{(1-\gamma)^{2}}{\gamma^{2}}t^{2}M_{k})\,.$ (42)
B.2.1 Estimation Error Bound for $\hat{Q}_{h}$
We start by defining some useful Bellman operators. The Bellman operator $J:{{\mathbb{R}}}^{S}\to{{\mathbb{R}}}^{S\times A}$ and its sampled version $\hat{J}:{{\mathbb{R}}}^{S}\to{{\mathbb{R}}}^{S\times A}$ are defined for every $V\in{{\mathbb{R}}}^{S}$ and any $(s,a)\in\mathcal{S}\times\mathcal{A}$ by
$JV(s,a)$ $:=r(s,a)+\gamma PV(s,a)\,,$ (43)
$\hat{J}V(s,a)$ $:=r(s,a)+\frac{\gamma}{m}\sum_{s^{\prime}\in\mathcal{C}(s,a)}V(s^{\prime})\,,$ (44)
where $\mathcal{C}(s,a)$ is a set of $m$ states sampled i.i.d from the distribution $P(\cdot|s,a)\,.$
Lemma B.6.
The operators $T$ and $\hat{T}$ satisfy for every $U,V\in{{\mathbb{R}}}^{S},s\in\mathcal{S},a\in\mathcal{A}$,
$|JU(s,a)-JV(s,a)|$ $\leq\gamma\lVert U-V\rVert_{\infty}\,,$ (45)
$|\hat{J}U(s,a)-\hat{J}V(s,a)|$ $\leq\gamma\max_{s^{\prime}\in\mathcal{C}(s,a)}|U(s)-V(s)|\,.$ (46)
Proof.
The proof is immediate and is hence omitted. ∎
We now show how to bound the error of $\hat{Q}_{k}$ for any $k\in\left[2,h\right]$. We can bound the total error on $\hat{Q}_{k}$ by decomposing the error into parts, accounting individually for the contribution of each stage in the estimation procedure. We will make use of the operators $\hat{J}$ as defined above and we will index them by the sets $\mathcal{S}_{k}$ ($k\in[1,h]$) for clarity. We use the notation $\|\cdot\|_{\mathcal{X}}$ for the infinity norm over any finite set $\mathcal{X}\,.$ Recall that $Q^{\pi}_{k}=r+\gamma P{\mathcal{T}}^{k-1}V^{\pi}=J{\mathcal{T}}^{k-1}V^{\pi}$ for every $k\in\left[1,h\right]$. The following decomposition holds for any $k\in\left[2,h\right]$ using the aforementioned contraction properties:
$\max_{a\in\mathcal{A}}\lVert\hat{Q}_{k}(\cdot,a)-Q^{\pi}_{k}(\cdot,a)\rVert_{\mathcal{S}_{k}}$
$=\max_{a\in\mathcal{A}}\lVert(\hat{J}_{\mathcal{S}_{k-1}}\hat{V}_{k})(\cdot,a)-(J{\mathcal{T}}^{k-1}V^{\pi})(\cdot,a)\rVert_{\mathcal{S}_{k}}$
$\leq\max_{a\in\mathcal{A}}\lVert(\hat{J}_{\mathcal{S}_{k-1}}\hat{V}_{k})(\cdot,a)-(\hat{J}_{\mathcal{S}_{k-1}}{\mathcal{T}}^{k-1}V^{\pi})(\cdot,a)\rVert_{\mathcal{S}_{k}}+\max_{a\in\mathcal{A}}\lVert(\hat{J}_{\mathcal{S}_{k-1}}{\mathcal{T}}^{k-1}V^{\pi})(\cdot,a)-(J{\mathcal{T}}^{k-1}V^{\pi})(\cdot,a)\rVert_{\mathcal{S}_{k}}$
$\leq\gamma\lVert\hat{V}_{k}-{\mathcal{T}}^{k-1}V^{\pi}\rVert_{\mathcal{S}_{k-1}}+\max_{a\in\mathcal{A}}\lVert(\hat{J}_{\mathcal{S}_{k-1}}{\mathcal{T}}^{k-1}V^{\pi})(\cdot,a)-(J{\mathcal{T}}^{k-1}V^{\pi})(\cdot,a)\rVert_{\mathcal{S}_{k}}$
$=\gamma\lVert M\hat{Q}_{k-1}-MJ{\mathcal{T}}^{k-2}V^{\pi}\rVert_{\mathcal{S}_{k-1}}+\max_{a\in\mathcal{A}}\lVert(\hat{J}_{\mathcal{S}_{k-1}}{\mathcal{T}}^{k-1}V^{\pi})(\cdot,a)-(J{\mathcal{T}}^{k-1}V^{\pi})(\cdot,a)\rVert_{\mathcal{S}_{k}}$
$\leq\gamma\max_{a\in\mathcal{A}}\lVert\hat{Q}_{k-1}(\cdot,a)-(J{\mathcal{T}}^{k-2}V^{\pi}(\cdot,a)\rVert_{\mathcal{S}_{k-1}}+\max_{a\in\mathcal{A}}\lVert(\hat{J}_{\mathcal{S}_{k-1}}{\mathcal{T}}^{k-1}V^{\pi})(\cdot,a)-(J{\mathcal{T}}^{k-1}V^{\pi})(\cdot,a)\rVert_{\mathcal{S}_{k}}$
$=\gamma\max_{a\in\mathcal{A}}\lVert\hat{Q}_{k-1}(\cdot,a)-Q^{\pi}_{k-1}(\cdot,a)\rVert_{\mathcal{S}_{k-1}}+\max_{a\in\mathcal{A}}\lVert(\hat{J}_{\mathcal{S}_{k-1}}{\mathcal{T}}^{k-1}V^{\pi})(\cdot,a)-(J{\mathcal{T}}^{k-1}V^{\pi})(\cdot,a)\rVert_{\mathcal{S}_{k}}$ (47)
which, after defining the terms $\Delta_{k}=\max_{a\in\mathcal{A}}\lVert\hat{Q}_{k}(\cdot,a)-Q^{\pi}_{k}(\cdot,a)\rVert_{\mathcal{S}_{k}}$ and
$\mathcal{E}_{k}=\max_{a\in\mathcal{A}}\lVert(\hat{J}_{\mathcal{S}_{k-1}}{\mathcal{T}}^{k-1}V^{\pi})(\cdot,a)-(J{\mathcal{T}}^{k-1}V^{\pi})(\cdot,a)\rVert_{\mathcal{S}_{k}}$ yields the following recursion:
$\Delta_{k}\leq\gamma\Delta_{k-1}+\mathcal{E}_{k}\,,$ (48)
and unrolling this recursion leads to the following bound on $\Delta_{h}$:
$\Delta_{h}\leq\gamma^{h-1}\Delta_{1}+\sum_{i=2}^{h}\gamma^{h-i}\mathcal{E}_{i}\,.$ (49)
We begin by bounding $\Delta_{1}$ by partially applying the recursion above:
$\Delta_{1}=\max_{a\in\mathcal{A}}\lVert\hat{Q}_{1}(\cdot,a)-JV^{\pi}(\cdot,a)\rVert_{\mathcal{S}_{1}}=\max_{a\in\mathcal{A}}\lVert\hat{J}_{\mathcal{S}_{0}}\hat{V}_{1}(\cdot,a)-JV^{\pi}(\cdot,a)\rVert_{\mathcal{S}_{1}}\leq\gamma\lVert\hat{V}_{1}-V^{\pi}\rVert_{\mathcal{S}_{0}}\,,$ (50)
which from Lemma B.5 we know to be bounded with probability at most $1-\delta|\mathcal{S}_{0}|$ by:
$\Delta_{1}\leq\gamma\lVert\hat{V}_{1}-V^{\pi}\rVert_{\mathcal{S}_{0}}\leq\frac{\gamma^{H+1}}{1-\gamma}+\frac{\gamma}{1-\gamma}\sqrt{\frac{\log(2|\mathcal{S}_{0}|/\delta)}{M_{0}}}\,.$ (51)
Finally, each term $\mathcal{E}_{j}$ for $j\leq k$ can be bounded using the following lemma below. The proof relies on a union bound over the states reached during the construction of the search tree, see the lemma towards the end of Szepesvári [2023] for a detailed explanation and a proof.
Lemma B.7.
[…]
$\max_{a\in\mathcal{A}}|(\hat{J}_{\mathcal{S}_{j-1}}{\mathcal{T}}^{k-1}V^{\pi})(S_{i},a)-(J{\mathcal{T}}^{k-1}V^{\pi})(S_{i},a)|\leq\frac{\gamma}{1-\gamma}\sqrt{\frac{\log(2/\delta)}{M_{j}}}\,.$ (52)
Combining this result with (51) and the unrolled recursion gives the final concentration bound with probability at least $1-\delta_{V}-\delta_{J}$ for $\delta_{V},\delta_{J}\in(0,1)\,,$
$\max_{a\in\mathcal{A}}\lVert\hat{Q}_{k}(\cdot,a)-Q^{\pi}_{k}(\cdot,a)\rVert_{\mathcal{S}_{k}}\leq\frac{\gamma^{H+k}}{1-\gamma}+\frac{\gamma^{k}}{1-\gamma}\sqrt{\frac{\log(2|\mathcal{S}_{0}|/\delta_{V})}{M_{0}}}+\frac{\gamma^{2}(1-\gamma^{k-1})}{(1-\gamma)^{2}}\sqrt{\frac{\log(2|\mathcal{A}||\mathcal{S}_{0}|(k-1)/\delta_{J})}{M}}$ (53)
by a union bound over $j\leq k$, assuming $M_{j}=M$ for all $11\leq j\leq h-$ 1.
B.2.2 End of the Proof of Theorem 5.4
For any state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, define the set $|\mathcal{S}_{h}|=\{s\}$ and the sets $|\mathcal{S}_{k}|$ accordingly for $10\leq k\leq h-$ 1. Here we set $M_{t}:=M>0\,,\forall t\in\left[1,h\right]$. Assuming the worst case growth of $|S_{k}|\leq|S_{k+1}||\mathcal{A}|M_{k+1}$, we can bound $|\mathcal{S}_{0}|$ deterministically as $|\mathcal{S}_{0}|\leq|\mathcal{A}|^{h}M_{h}M^{h-1}$. Using the bound in the previous section, it is possible to bound the error of $\hat{Q}_{h}(s,a)$ as follows:
$|\hat{Q}_{h}(s,a)-Q_{h}^{\pi}(s,a)|$ $\leq\max_{a^{\prime}\in\mathcal{A}}\lVert\hat{Q}_{h}(\cdot,a^{\prime})-Q^{\pi}_{h}(\cdot,a^{\prime})\rVert_{\mathcal{S}_{h}}$
$\leq\frac{\gamma^{H+h}}{1-\gamma}+\frac{\gamma^{h}}{1-\gamma}\sqrt{\frac{\log(2|\mathcal{S}_{0}|/\delta_{V})}{M_{0}}}+\frac{\gamma^{2}(1-\gamma^{h-1})}{(1-\gamma)^{2}}\sqrt{\frac{\log(2|\mathcal{A}||\mathcal{S}_{0}|h/\delta_{J})}{M}}\,,$ (54)
with probability at least $1-\delta_{V}-\delta_{J}$.
Assume now we want to estimate the value function $\hat{Q}(s,a)$ for all state action pairs $(s,a)$ in some set $\mathcal{C}\subseteq\mathcal{S}\times\mathcal{A}$: we want to achieve $\max_{(s,a)\in\mathcal{C}}|\hat{Q}(s,a)-Q(s,a)|\leq b$ for some positive accuracy $b$ with probability at least $1-\delta_{V}-\delta_{J}$. This can be achieved by choosing the following parameters:
$\forall 1\leq j\leq h\,,M_{j}:=M$ $>\frac{9\gamma^{4}(1-\gamma^{h-1})^{2}}{(1-\gamma)^{4}b^{2}}\log(2h|\mathcal{A}||\mathcal{S}_{0}||\mathcal{C}|/\delta_{J})\,,$ (55)
$M_{0}$ $>\frac{9\gamma^{2h}}{(1-\gamma)^{2}b^{2}}\log(2|\mathcal{S}_{0}||\mathcal{C}|/\delta_{V})\,,$ (56)
$H$ $>\frac{1}{1-\gamma}\log\left(\frac{3\gamma^{h}}{b(1-\gamma)}\right)\,.$ (57)
Picking $\mathcal{C}=\mathcal{S}\times\mathcal{A}$ gives $\lVert\hat{Q}_{h}-Q^{\pi}_{h}\rVert_{\infty}\leq b$. If estimates for $\{V_{k}\}_{k=1}^{h}$ are reused across state action pairs, and the components of each value function are only computed at most once for every state $s\in|\mathcal{S}|$, then the overall number of samples is of the order of $KM_{0}H|\mathcal{S}_{0}|+Kh|\mathcal{S}_{0}|\cdot|\mathcal{C}|M$. In particular, with $\mathcal{C}=\mathcal{S}\times\mathcal{A}$, the number of samples becomes of the order of $KM_{0}H|\mathcal{S}|+Kh|\mathcal{S}|\cdot|\mathcal{A}|M$ where we reuse estimators (which are in this case computed for each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$. Let $K>\frac{1}{h(1-\gamma)}\log(\frac{4}{\epsilon(1-\gamma)(1-\gamma^{h})})$ as in Theorem 5.4. Choosing $\delta_{V}=\delta_{J}=\delta/2$ for some $\delta\in(0,1)$ and $b=\frac{\epsilon(1-\gamma)(1-\gamma^{h})}{4}$ yields the following bound with probability at least $1-\delta$ (by Theorem 5.3):
$\lVert V^{\star}-{V^{\pi_{k}}}\rVert_{\infty}\leq\gamma^{hk}\left(\lVert V^{\star}-V^{\pi_{0}}\rVert_{\infty}+\frac{1-\gamma^{hk}}{(1-\gamma)(1-\gamma^{h})}\right)+\frac{\epsilon}{2}\,.$ (58)
To conclude, note that when $k=K$, we have:
$\gamma^{hK}\left(\lVert V^{\star}-V^{\pi_{0}}\rVert_{\infty}+\frac{1-\gamma^{hK}}{(1-\gamma)(1-\gamma^{h})}\right)\leq\gamma^{hK}\frac{2(1-\gamma^{hK})}{(1-\gamma)(1-\gamma^{h})}\leq\frac{\epsilon(1-\gamma^{hK})}{2}\leq\frac{\epsilon}{2}\,.$ (59)
Appendix C Additional Details and Proofs for Section 6: $h$-PMD with Function Approximation
To incorporate function approximation to $h$-PMD, the overall approach is as follows: compute an estimate of the lookahead value function by fitting a linear estimator using least squares, then perform an update step as in inexact PMD, taking advantage of the same error propagation analysis which is agnostic to the source of errors in the value function. However a problem that arises is in storing the new policy, since a classic tabular storage method would require $\mathcal{O}(|\mathcal{S}|)$ memory, which we would like to avoid to scale to large state spaces. A solution to this problem will be presented later on in this section.
Since we are using lookahead in our $h$-PMD algorithm, we modify the targets in our least squares regression procedure in order to approximate $Q_{h}^{\pi}$ rather than $Q^{\pi}$ as in the standard PMD algorithm (with $h=1$).
C.1 Inexact $h$-PMD with Linear Function Approximation
Here we provide a detailed description of the $h$-PMD algorithm using function approximation to estimate the $Q_{h}$-function. Firstly we consider the lookahead depth $h$ as a fixed input to the algorithm. We assume that the underlying MDP has finite state and action spaces $\mathcal{S}$ and $\mathcal{A}$, with transition kernel $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{A})$, reward vector $R\in\left[0,1\right]^{|\mathcal{S}|\times|\mathcal{A}|}$ and discount factor $\gamma$. We also assume we are given a feature map $\psi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d}$, also written as $\Psi\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times d}$.
We adapt the inexact $h$-PMD algorithm from section 5 as follows:
  1. 1.
    Take as input any policy $\pi_{0}\in\text{rint}(\Pi)$, for example one that is uniform over actions in all states.
  2. 2.
    Initialize our memory containing $\pi_{0}$, and an empty list $\Theta$ which will hold the parameters of our value function estimates computed at each iteration.
  3. 3.
    At each iteration $k=0,\ldots,K$, perform the following steps:
    1. (a)
      Compute the targets $\hat{R}(z):=\hat{Q}_{h}(z)$ for all $z\in\mathcal{C}$ using the procedure described in section 5.1, using $\pi_{k}$ as a policy.
    2. (b)
      Compute $\theta_{k}:=\operatorname{argmin}_{\theta\in\mathbb{R}^{d}}\sum_{z\in\mathcal{C}}\rho(z)(\psi(z)^{\top}\theta-\hat{R}(z))^{2}$ and append it to the list $\Theta$.
    3. (c)
      Using $\Psi\theta_{k}$ as an estimate for $Q^{\pi_{k}}_{h}$, compute ${\pi_{k+1}}$ using the usual update from inexact $h$-PMD. This step can be merged with step 3a, the details are described below.
Computing the full policy $\pi_{k}$ for every state at every iteration would require $\Omega(|\mathcal{S}|)$ operations, although the number of samples to the environment still depends only on $d$. If this is not a limitation the policy update step can be performed as in inexact $h$-PMD simply by using $\Psi\theta_{k}$ as an estimate for $Q_{h}^{\pi_{k}}$, as described in step 3c. Otherwise, we describe a method below to compute $\pi_{k}$ on demand when it is needed in the computation of the targets in step 3a, eliminating step 3c entirely.
  1. 1.
    The policy $\pi_{0}$ is already given for all states.
  2. 2.
    For any $k>0$ given ${\pi_{k}}$ we can compute ${\pi_{k+1}}$ at any state $s$ using the following:
    1. (a)
      Compute any greedy policy $\tilde{\pi}(\cdot|s)$ using $\Psi\theta_{k}$, for example $\tilde{\pi}(a^{\prime}|s)\propto\mathbbm{1}\left\{a^{\prime}\in\underset{a\in\mathcal{A}}{\operatorname{argmax}}\left\{\langle\psi(s,a)^{\top}\theta_{k}\right\}\right\}$
    2. (b)
      Compute $\eta_{k}:=\frac{1}{\gamma^{2h(k+1)}}D_{\phi}(\tilde{\pi}(\cdot|s),\pi^{k}(\cdot|s))$
    3. (c)
      Finally, compute $\pi_{k+1}(\cdot|s)\in\operatorname{argmax}_{\pi\in\Pi}\left\{\eta_{k}\langle(\Psi\theta_{k})_{s},\pi(\cdot|s)\rangle-D_{\phi}(\pi(\cdot|s),\pi_{k}(\cdot|s))\right\}$
Remark C.1.
This procedure can be used to compute ${\pi_{k+1}}(\cdot|s)$ “on demand” for any $k>0$ and $s\in\mathcal{S}$ by recursively computing ${\pi_{k}}(\cdot|s^{\prime})$ at any states $s^{\prime}$ required for the computation of $\Psi\theta_{k}$ (since $\pi_{0}$ is given). The memory cost requirement for this method is of the order of $\mathcal{O}(K|\mathcal{A}|\min\{|\mathcal{S}_{0}|H,|\mathcal{S}|\})$ in total which is always less than computing the full tabular policy at each iteration.
Remark C.2.
Above, $(\Psi\theta_{k})_{s}$ denotes a vector in $\mathbb{R}^{|\mathcal{A}|}$ that can be computed in $\mathcal{O}(|\mathcal{A}|d)$ operations: each component in $a$ is equal to $\psi(s,a)^{\top}\theta_{k}$ which can be computed in $\mathcal{O}(d)$ operations. Also, the stepsize $\eta_{k}$ is chosen to be equal to its lower bound in the analysis of 5.3, with the value of $c_{k}$ chosen to be equal to $\gamma^{2h(k+1)}$, as it is needed for linear convergence. We also note that by memorizing policies as they are computed, we manage to perform just as many policy updates as are needed by the algorithm.
C.2 Convergence Analysis of Inexact $h$-PMD using Function Approximation
Our analysis follows the same lines as the proof in [Lattimore and Szepesvari, 2019] under our assumptions from section 6.
We compute $\hat{Q}_{h}(z)$ for a subset $\mathcal{C}$ of state action pairs using the vanilla Monte Carlo Planning algorithm presented in section 5. These will be our targets for estimating the value function using least squares. In order to control the extrapolation of these estimates to the full $Q_{h}^{\pi}$ function, we will need the following theorem.
Theorem C.3 ([Kiefer and Wolfowitz, 1960]).
Assume $\mathcal{Z}$ is finite and $\psi:\mathcal{Z}\to\mathbb{R}^{d}$ is such that the underlying matrix $\Psi\in\mathbb{R}^{|\mathcal{Z}|\times d}$ is of rank $d$. Then there exist $\mathcal{C}\subseteq\mathcal{Z}$ and a distribution $\rho:\mathcal{C}\to\left[0,1\right]$ (i.e. $\sum_{z\in\mathcal{C}}\rho(z)=1$) satisfying the following conditions:
  1. 1.
    $|\mathcal{C}|\leq d(d+1)/2$ ,
  2. 2.
    $\sup_{z\in\mathcal{Z}}\lVert\psi(z)\rVert_{\mathcal{G}_{\rho}^{-1}}\leq\sqrt{d}$ where $\mathcal{G}_{\rho}:=\sum_{z\in\mathcal{C}}\rho(z)\psi(z)\psi(z)^{\top}$.
Let $\mathcal{C}\subseteq\mathcal{S}\times\mathcal{A}$, $\rho:\mathcal{C}\to\left[0,1\right]$ be as described in the Kiefer-Wolfowitz theorem. Define $\hat{\theta}$ as follows:
$\hat{\theta}\in\operatorname{argmin}_{\theta\in\mathbb{R}^{d}}\sum_{z\in\mathcal{C}}\rho(z)(\phi(z)^{\top}\theta-\hat{R}_{m}(z))^{2}\,.$ (60)
Assuming that $\mathcal{G}_{\rho}$ is nonsingular, a closed form solution for the above least squares problem is given by $\hat{\theta}=\mathcal{G}_{\rho}^{-1}\sum_{z\in\mathcal{C}}\rho(z)\hat{R}_{m}(z)\phi(z)$. This will be used to estimate our lookahead value function as $\Psi\hat{\theta}$. We can control the extrapolation error of this estimate as follows:
Theorem C.4.
Let $\delta\in(0,1)$ and let $\theta\in\mathbb{R}^{d}$ such that $\lVert Q_{h}^{\pi}-\Phi\theta\rVert_{\infty}\leq\epsilon_{\pi}$. Assume that we estimate the targets $\hat{Q}_{h}$ using the Monte Carlo planning procedure described in section 5.1 with parameters, $M$ and $M_{0}$. There exists a subset $\mathcal{C}\subseteq\mathcal{Z}$ and a distribution $\rho:\mathcal{C}\to\mathbb{R}_{\geq 0}$ for which, with probability at least $1-\delta$, the least squares estimator $\hat{\theta}$ satisfies:
$\lVert Q_{h}^{\pi}-\Phi\hat{\theta}\rVert_{\infty}\leq\epsilon_{\pi}(1+\sqrt{d})+\epsilon_{Q}\sqrt{d}\,,$ (61)
where
$\epsilon_{Q}:=\frac{\gamma^{H+h}}{1-\gamma}+\frac{\gamma^{h}}{1-\gamma}\sqrt{\frac{\log(4Zd^{2}/\delta)}{M_{0}}}+\frac{\gamma^{2}(1-\gamma^{h-1})}{(1-\gamma)^{2}}\sqrt{\frac{\log(4hZd^{2}/\delta)}{M}}\,,$ (62)
and Z := $|\mathcal{A}|^{h}M_{h}M^{h-1}\,.$
Proof.
The proof of this theorem is identical to the one in Lattimore and Szepesvari [2019], up to using our new targets and error propagation analysis instead of the typical function approximation targets for $Q$-function estimation (without lookahead). The subset $\mathcal{C}\subseteq\mathcal{S}\times\mathcal{A}$ and distribution $\rho:\mathcal{C}\to\left[0,1\right]$ are given by the Kiefer-Wolfowitz theorem stated above. We suppose we have access to such a set. Note that this assumption is key to controlling the extrapolation error of a least squares estimate $\Psi\hat{\theta}$. ∎
We have now fully described how to estimate the lookahead value function $Q^{\pi}_{h}$ using a number of trajectories that does not depend on the size of the state space. In the previous section, we described how to use this estimate in an $h$-PMD style update, while maintaining a memory size that does not depend on $|\mathcal{S}|$. Therefore we have an $h$-PMD algorithm that does not use memory, computation or number of samples depending on $|\mathcal{S}|$.
Appendix D Additional Details and Simulations
Here we provide details of an additional experiment running $h$-PMD with Euclidean Divergence. As in section 7, experiments were run on the “bsuite” DeepSea environment with $\gamma=0.99$ and a grid size of 64x64. The figures are identical in structure to the ones in 7, 32 runs were performed for each value of $h$ in 1, 5, 10, 15, and 20. In Fig. 3, we report the number of samples to the environment at each iteration for the same runs. When compared to the figures showing the convergence, it is clearly seen that running the algorithm with higher values of lookahead results in convergence using fewer overall samples.
Figure 2: Suboptimality value function gap for $h$-PMD using Euclidean Bregman divergence. in the exact (left) and inexact (right) settings. (Right) We performed 32 runs for each value of $h$, the mean is shown as a solid line and the standard deviation as a shaded area.
Figure 3: Samples used by inexact $h$-PMD at each iteration. When compared with the first figure, it is clear to see the algorithm needs less samples to converge with higher values of $h$. Since far less iterations are needed with higher values of lookahead, the benefit of higher convergence rate greatly outweighs the additional cost per iteration.